### v1

# Multimodal Graph Benchmark

Abstract
Associating unstructured data with structured information is crucial for real-world tasks that require relevance search. However, existing graph learning benchmarks often overlook the rich semantic information associate with each node. To bridge such gap, we introduce the Multimodal Graph Benchmark (MM-Graph), the first comprehensive multi-modal graph benchmark that incorporates both textual and visual information. MM-Graph surpasses previous efforts, which have primarily focused on text-attributed graphs with various connectivity patterns. MM-Graph consists of five graph learning datasets of various scales that are appropriate for different learning tasks. Their multimodal node features, enabling a more comprehensive evaluation of graph learning algorithms in real-world scenarios. To facilitate research on multimodal graph learning, we further provide an extensive study on the performance of various graph neural networks in the presence of features from various modalities. MM-Graph aims to foster research on multimodal graph learning and drive the development of more advanced and robust graph learning algorithms. By providing a diverse set of datasets and benchmarks, MM-Graph enables researchers to evaluate and compare their models in realistic settings, ultimately leading to improved performance on real-world applications that rely on multimodal graph data.
1 Introduction
Graphs are ubiquitous data structures that can effectively represent complex relationships and interactions between entities in various domains, such as social networks, biological systems, and recommendation systems. In real-world scenarios, these entities often possess rich semantic information in the form of unstructured data, such as images and text descriptions. Associating these unstructured data with the structured graph information is essential for tasks that require relevance search and information retrieval, such as recommendation systems [27, 3, 10].
Despite the importance of multimodal graph learning, existing graph learning benchmarks typically focus on investigating graphs with various connectivity patterns, such as long-range links, out-of-distribution scenarios, etc [5, 9, 19, 20, 13]. These benchmarks aim to evaluate the performance of graph neural networks (GNNs) and other graph learning algorithms on different graph structures and topologies. However, these benchmarks under-utilize the rich semantic information naturally present in each node, limiting their ability to evaluate graph learning algorithms in realistic settings.
With the prevalence of language modeling, there has been an outbreak in the effective integration of text and graph topology on text-attributed graphs, leading to a comprehensive benchmark on text-attributed graphs [33, 35, 15, 23]. While textual information undoubtedly plays a crucial role in understanding the data and making predictions for user-centric tasks, it is essential to recognize that other modalities, particularly visual information, can be helpful and provide complementary information that text alone cannot capture. For example, as shown in Figure  1(a), visually similar products may have distinct text features and ignoring such semantic gap between modalities may prevent GNNs from fully exploiting the relationships between multimodal semantics and structures.
Figure 1: Visualization of our Multimodal Graph Benchmark. All nodes of our benchmark have both visual and text features. (a) Amazon-Sports: The image and text come from the original image and title of the sports equipment. (b) Goodreads-LP: The image comes from the cover of the book. We do not show the text features of Goodreads-LP since the book description is very long. (c) Ele-fashion: The image and text come from the original image and title of the fashion product.
To address this limitation, we introduce the Multimodal Graph Benchmark (MM-Graph), the first comprehensive graph benchmark that incorporates both textual and visual information. MM-Graph consists of five graph learning datasets of various scales, designed for diverse downstream tasks such as node classification and link prediction. These datasets contain multimodal node features, including text and images, alongside the traditional graph structure. By providing a diverse set of datasets and benchmarks, MM-Graph enables researchers to evaluate and compare their models in realistic settings, fostering research on multimodal graph learning. MM-Graph also standardizes GNN architectures, feature encoders, and dataloaders and evaluators in order to enable a comprehensive evaluation on multimodal graph data
To show the importance of these datasets in for advancing multimodal graph learning , we adapt two state-of-the-art multimodal graph neural networks from the recommendation systems to the graph learning domain. These models have demonstrated strong performance in leveraging multimodal information for recommendation tasks. By applying them to the datasets in MM-Graph, we aim to provide a comprehensive evaluation of their effectiveness in various graph learning scenarios. We conduct an extensive study on the performance of MMGCN [31], MGAT [28], and other popular graph neural network models, such as GraphSAGE [11], in the presence of multimodal node features.
Surprisingly, we find that GNNs specifically designed for multimodal scenarios, such as MMGCN and MGAT, do not consistently outperform general-purpose GNN models like GraphSAGE across all datasets in MM-Graph. This finding highlights the need for a standard multimodal graph benchmark to rigorously evaluate and compare the performance of different models. MM-Graph fills this gap by providing a diverse set of datasets with varying scales, tasks, and modality combinations. By establishing a common ground for evaluation, MM-Graph enables researchers to gain a deeper understanding of the strengths and limitations of existing multimodal graph learning algorithms and identify areas for improvement. Moreover, MM-Graph opens up new research opportunities in multimodal graph learning. The availability of datasets with rich textual and visual information allows researchers to explore novel architectures and techniques that can effectively capture and integrate the semantic information from multiple modalities. This can lead to the development of more advanced and robust graph learning algorithms that can handle the complexities of real-world multimodal graph data.
Contributions and novelty. In this work, we introduce a multimodal graph benchmark known as MM-Graph. As design principles, we strive to create non-trivial, realistic data that comes from real-world applications and will benefit real-world deployments. Specifically, MM-Graph contains three link prediction datasets and two node classification datasets, as shown in Table  1. We summarize our novel contributions as follows: (1) To the best of our knowledge, no existing standardized graph benchmark provides both text and image features. This is important as images provide complementary context and details. While one can extract keywords and intent from text, images supply visual cues about objects, scenes, emotions, etc. that give richer understanding, especially for more abstract or complex tasks. (2) We provide a comprehensive analysis on the performance of various GNNs on MM-Graph, highlighting the need of developing new multimodal GNNs. (3) In terms of graph benchmarks, MM-Graph is the first benchmark to include not only various GNNs but also different ways of encoding node features, which presents an important and practical question of which feature encoder gives the best features for structure learning tasks.
Reproducibility: MM-Graph code, datasets and details are publicly available. The datasets are hosted on Huggingface. The evaluation code and baselines are hosted on GitHub with documentations.
Table 1: Overview of the datasets in the proposed MM-Graph. We present five multimodal graph datasets with varying scale and tasks. All datasets are splitted in random. LP = Link Prediction. NC = Node Classification.
Name Text Features Visual Features Proposed Task Metrics Scale Split Ratio
Amazon-Sports Product Titles Product Images LP MRR, Hits@K Small 8/1/1
Amazon-Cloth Product Titles Product Images LP MRR, Hits@K Medium 8/1/1
Goodreads-LP Book Description Book Images LP MRR, Hits@K Large 6/1/3
Ele-fashion Fashion Titles Fashion Images NC Accuracy Medium 6/1/3
Goodreads-NC Book Description Book Images NC Accuracy Large 6/1/3
2 Related Work
Multimodal Feature Learning. Multimodal feature learning aims to learn joint representations from multiple modalities, such as text, images, and audio. Transformer-based models have shown remarkable success in multimodal feature learning  [32, 24]. These models learn typically transferable visual representations by leveraging corresponding natural language supervision. Models like FLAVA  [26] and Perceiver [14] have demonstrated the effectiveness of jointly pre-training transformers on unpaired images and text while CLIP has shown that contrastive objectives can effectively align representations from different modalities  [26, 14, 24]. We refer to the survey for more details  [32].
Multimodal Graph Learning. While most existing multimodal learning approaches focus on 1-to-1 mappings between modalities, real-world data often exhibits more complex many-to-many relationships that can be represented as graphs. Multimodal graph learning (MMGL) aims to address this challenge by leveraging graph structures to learn from multimodal data with complicated relations among multiple multimodal neighbors  [6, 34]. Recent efforts in MMGL have focused on integrating GNNs with pretrained language models (LMs) [33]. Our work builds upon the recent advances in multimodal feature learning and MMGL by introducing a comprehensive benchmark that incorporates both structured and unstructured modalities. We adapt up-to-date multimodal feature learning frameworks and state-of-the-art multimodal GNNs and provide an extensive study on their performance in the presence of features from various modalities.
Benchmarks for graph representation learning. Several established graph benchmarks have been developed and widely adopted [20, 13, 7]. However, when it comes to learning on graphs with rich features, these benchmarks exhibit notable deficiencies. Firstly, a majority of these datasets suffer from the absence of raw feature information, limiting the investigation of attribute modeling’s effectiveness. Secondly, these datasets often neglect to explore the impact of feature modeling on GNNs. Thus, there is a compelling necessity to construct a comprehensive graph benchmark with rich multimodal features. Yan et al. proposed the first standardized graph benchmark with rich textual information, fostering the development of encompassing GNNs with prevalent language models [33]. However, we argue that it is critical to investigate modalities other than text, especially images in structure learning tasks and proposing the first standarized multimodal graph benchmark to foster the development of research in this direction.
3 MM-Graph: The Multimodal Graph Benchmark
In this section, we discuss the proposed 5 real-world datasets in MM-Graph. The basic information is shown in Table  1. We first provide diverse graph-related statistics so as to give a comprehensive understanding of the structural properties of the datasets. This information can guide the selection of appropriate graph learning algorithms and help interpret the performance of different models on each dataset.
3.1 Data Curation
We first introduce the data curation process of all datasets, as long as their license.
Amazon-Sports Amazon-Sports is a link prediction dataset, based on the Amazon-Review dataset [21, 12], where each node corresponds to a product on Amazon in the sports category and the link captures whether two products are co-purchased together. The text features are the titles of the products and the visual features are the high-resolution raw images of the products. In Figure  1(a), we show an example of nodes and edges in Amazon-Sports. The visual features are images of tents of various shapes under diverse backgrounds and text features are the title of the tents. Specifically, we use the data pre-processing script provided by Patton [16].
Amazon-Cloth Similar to Amazon-Sports, Amazon-Cloth is constructed according to the co-purchasing information of Amazon products in the clothes category. The text features are titles of products, e.g. Nike mens Revolution 6 Road Running and the visual features are corresponding images. We follow the same data curation procedure.
Goodreads-LP is based on the Goodreads Book Graph dataset  [29, 30]. Here we construct the graph in the way where each node corresponds to a book on Goodreads and the link captures if a user who likes this book will like the other book, following  [35]. The text features are the descriptions of the books and the visual features are the cover of the book. Nodes without images are removed.
Goodreads-NC is a node classification dataset, based on the Goodreads dataset [29, 30], where each node corresponds to a book on Goodreads and the link captures if a user who likes this book will like the other book. The text features are the descriptions of the books and the visual features are the cover image of the book. Nodes without images are removed.
Ele-fashion Ele-fashion is a node classification dataset, based on the Amazon-Fashion dataset [21, 12], where each node corresponds to a product and the link captures if a user who buy this product will also buy the other product. The text features are the titles of the products and the visual features are the images of the products.
Data Availability and Ethics. Our benchmark is organized from existing open source data, with proper open source licenses. Amazon-Sports and Amazon-Cloth are available with Apache License  [^1] https://github.com/PeterGriffinJin/Patton. Goodreads-LP, Goodreads-NC and Ele-fashion are released under MIT License  [^2] https://mengtingwan.github.io/data/goodreads.html. These do not involve interaction with humans or private data.
Table 2: Detailed graph-based statistics of each dataset proposed in MM-Graph. CC = Cluster Coefficient. RA = Resource Allocation. N/A = Not Applicable (Nodes do not have node labels). Detailed definition for each statistics provided in the Appendix.
Name Nodes Edges Average Degree Average CC Average RA Transitivity Edge Homophily
Amazon-Sports 50,250 356,202 14.18 0.4002 0.3377 0.2658 N/A
Amazon-Cloth 125,839 951,271 15.12 0.2940 0.2588 0.1846 N/A
Goodreads-LP 636,502 3,437,017 10.79 0.1102 0.0685 0.0348 N/A
Goodreads-NC 685,294 7,235,084 21.11 0.1614 0.1056 0.0498 0.6667
Ele-fashion 97,766 199,602 4.08 0.1730 0.1467 0.0560 0.7675
3.2 Tasks
Link Prediction. The task is to predict new association edges given the training edges. The evaluation is based on how well a model ranks positive test edges over negative test edges. Specifically, for Amazon-Sports and Amazon-Cloth, we generate hard valid/test negatives using HeaRT  [18]. HeaRT is recognized as a better way of generating negatives for link prediction tasks in the community. For Goodreads-LP, since it takes more than 120 hours to generate hard negatives using HeaRT, we perform random sampling. Each positive edge in the validation/test set is ranked against 150 hard negative edges. For evaluation metrics, we report MRR, Hits@10, and Hits@1 - the three most commonly-used evaluation metrics for link prediction [13, 18]. For Amazon-Sports and Amazon-Cloth, edges are randomly split into train/valid/test splits according to 8/1/1 ratio. For Goodreads-LP, edges are randomly split according to 6/1/3 ratio. Validation and test edges are explicitly removed from the graphs to avoid any potential leakage [36].
Node Classification. The task is to predict the book category from $N$ available categories. We report accuracy following  [13, 33]. For Goodreads-NC, there are 10 categories such as History, Children and comics. For Ele-Fashion, there are 12 categories such as shoes, jewelry and dresses. For evaluation metrics, we report accuracy, as it’s the most common metric for node classification[13]. Nodes are randomly split into train/valid/test following 6/1/3 ratio.
3.3 Statistics
We provide graph-related statistics are presented in Table  2. The detailed descriptions for each metric is provided in the Appendix. These statistics provide valuable insights into the structural properties of the graphs in each dataset.
Amazon-Sports.There are 50,250 nodes with a total of 356,202 edges in Amazon-Sports, which is a small-scale link prediction dataset. The small scale makes Amazon-Sports great for trying it at first, as it is computation and memory efficient for most applications. The average degree is 14, which indicates that each node have 14 neighbors on average.
Amazon-Cloth.There are 125,839 nodes with a total of 951,271 edges in Amazon-Cloth. Amazon-Cloth constitutes a medium-scale link prediction dataset. Since the scale of this dataset is much larger than Amazon-Sports, Amazon-Cloth serves as a solid benchmark for real-world e-commerce based applications. The average degree is 15, which similar to Amazon-Sports. But the average clustering coefficient is much smaller. This indicates that nodes in Amazon-Cloth are more likely to scatter around, instead of forming communities despite its high average degree.
Goodreads-LP.There are 636,502 nodes with a total of 3,437,017 edges, which constitutes a large-scale link prediction dataset. The average degree is 10.
Goodreads-NC. There are 685,294 nodes with a total of 7,235,084 edges, which constitutes a large-scale node classification dataset. The average degree is 21, highest among all datasets.
Ele-fashion. There are 97,766 nodes with a total of 199,602 edges, which constitutes a medium-scale node classification dataset. The average degree is 4, which is the most sparse dataset among MM-Graph.
4 Evaluation Framework
To enable a comprehensive evaluation on multimodal graph data, MM-Graph standardizes the GNN architectures, feature encoders, and dataloaders and evaluators used across all datasets.
4.1 Graph Neural Networks
To enable a comprehensive evaluation of GNN architectures on multimodal graph data, MM-Graph includes five representative GNN models: GCN [17], SAGE [11], MMGCN [31], MGAT [28], and BUDDY [2] . Additionally, following  [13], we report the performance of an MLP as a baseline to evaluation the usefulness of graph structure.
Conventional GNNs. We standardizes GCN, SAGE, MLP for both link prediction and node classification. Since BUDDY is specifically designed for link prediction, so we do not report its performance on node classification.
Multimodal GNNs. While GNNs have shown strong performance on tasks involving graphs with unimodal features, there has been limited work on developing GNNs that can effectively handle multimodal graph data. In order to evaluate the effectiveness of GNNs on multimodal graphs, we adapt two recent architectures from the recommendation systems domain to our standardized multimodal graph benchmarks:
MMGCN is a GCN-based framework that leverages information interchange between users and micro-videos in multiple modalities to refine their modality-specific representations and model users’ fine-grained preferences. It constructs separate user-item graphs for each modality. The information interchange of users and items in each modality is encoded using GNNs. And modality representations are fused in the prediction layer to predict possible recommendations [31]. For LP, we use the most common dot product decoder to decode possible links. For NC, we stack a 3-layer MLP as transform representations to a number of dimensions equal to the number of node classes.
MGAT extends the popular GAT architecture to multimodal settings. It learns modality-specific node representations which are then combined using a cross-modal attention layer  [28]. This allows MGAT to weigh different modalities based on their importance. Similar to MMGCN, we decode links using dot product for LP and perform NC through a 3-layer MLP.
4.2 Feature Encoders
Text encoders. For text encoders, we select CLIP  [24], T5  [25] and ImageBind  [8]. T5 is selected as the STOA text embedding models, while CLIP and ImageBind are selected so that the output text representations aligns with visual representations.
Visual encoders. For visual encoders, we select CLIP  [24], ViT  [4], ImageBind  [8] and Dinov2  [22]. ViT and Dinov2 are selected as two different types of visual encoders, while ViT are explicitly trained with supervision, Dinov2 learns robust visual features without supervision. CLIP and ImageBind are selected so that the output text representations aligns with visual representations. Specifically, Imagebind has potentials to extend multimodal graph learning to other modalities, such as audio and video thanks to its ability to bind all modalities into one embedding space. Studies on various feature encoders give insights on design choices of multimodal graph learning, such as (1) how important it is to align multiple modalities into a unified embedding space, (2) if supervised feature encoders work better than unsupervised feature encoders. Note that feature encoders are only used to extract features for GNNs and are never trained upon.
Table 3: Link prediction results on Amazon-Sports, Amazon-Cloth and Goodreads-LP. *OOM = out of GPU memory. Highlighted box indicates the best performing combination for each dataset.
Encoder Amazon-Sports Amazon-Cloth Goodreads-LP
Image Text MRR $\uparrow$ H@1 $\uparrow$ H@10 $\uparrow$ MRR $\uparrow$ H@1 $\uparrow$ H@10 $\uparrow$ MRR $\uparrow$ H@1 $\uparrow$ H@10 $\uparrow$
MMGCN CLIP 31.96 $\pm$0.10 16.35 $\pm$0.11 68.46 $\pm$0.08 22.20 $\pm$0.05 10.76 $\pm$0.1 46.62 $\pm$0.12 OOM OOM OOM
ViT T5 30.33 $\pm$0.03 15.01 $\pm$0.05 66.41 $\pm$0.11 19.45 $\pm$0.34 9.22 $\pm$0.20 40.49 $\pm$0.61 OOM OOM OOM
ImageBind 31.74 $\pm$0.21 16.45 $\pm$0.13 67.39 $\pm$0.74 24.72 $\pm$0.19 12.47 $\pm$0.09 51.32 $\pm$0.56 OOM OOM OOM
Dinov2 T5 30.04 $\pm$0.27 14.98 $\pm$0.07 64.56 $\pm$0.56 21.77 $\pm$0.23 10.47 $\pm$0.12 45.81 $\pm$0.52 OOM OOM OOM
MGAT CLIP 27.56 $\pm$0.30 13.55 $\pm$0.29 60.21 $\pm$0.21 21.38 $\pm$0.23 10.39 $\pm$0.22 44.60 $\pm$0.36 OOM OOM OOM
ViT T5 30.15 $\pm$0.34 15.28 $\pm$0.34 64.84 $\pm$0.41 20.59 $\pm$0.41 9.79 $\pm$0.30 43.44 $\pm$0.76 OOM OOM OOM
ImageBind 30.15 $\pm$0.12 15.50 $\pm$0.05 64.20 $\pm$0.43 22.13 $\pm$0.27 10.96 $\pm$0.15 45.84 $\pm$0.57 OOM OOM OOM
Dinov2 T5 28.91 $\pm$0.09 14.47 $\pm$0.18 62.11 $\pm$0.22 21.42 $\pm$0.13 10.38 $\pm$0.13 44.11 $\pm$0.50 OOM OOM OOM
GCN CLIP 31.38 $\pm$0.08 16.58 $\pm$0.13 66.14 $\pm$0.08 22.28 $\pm$0.05 11.83 $\pm$0.04 43.52 $\pm$0.10 25.34 $\pm$0.06 13.81 $\pm$0.12 50.36 $\pm$0.14
ViT T5 30.83 $\pm$0.07 16.31 $\pm$0.08 64.76 $\pm$0.15 21.60 $\pm$0.05 11.37 $\pm$0.03 42.29 $\pm$0.14 26.50 $\pm$0.10 14.86 $\pm$0.08 51.54 $\pm$0.14
ImageBind 31.67 $\pm$0.09 17.07 $\pm$0.14 65.61 $\pm$0.10 22.81 $\pm$0.03 12.27 $\pm$0.05 44.28 $\pm$0.09 27.56 $\pm$1.26 14.31 $\pm$1.37 57.25 $\pm$0.52
Dinov2 T5 30.42 $\pm$0.02 16.02 $\pm$0.03 64.02 $\pm$0.06 21.19 $\pm$0.08 11.09 $\pm$0.06 41.46 $\pm$0.16 28.21 $\pm$1.12 15.11 $\pm$1.06 57.94 $\pm$0.95
SAGE CLIP 33.83 $\pm$0.08 17.57 $\pm$0.14 71.90 $\pm$0.07 24.58 $\pm$0.18 12.16 $\pm$0.11 51.12 $\pm$0.09 44.10 $\pm$1.37 32.32 $\pm$1.38 69.07 $\pm$1.19
ViT T5 32.01 $\pm$0.10 15.94 $\pm$0.17 69.84 $\pm$0.21 23.11 $\pm$0.05 11.10 $\pm$0.04 48.89 $\pm$0.09 44.79 $\pm$0.18 33.11 $\pm$0.21 69.43 $\pm$0.18
ImageBind 34.32 $\pm$0.11 17.87 $\pm$0.23 73.04 $\pm$0.15 25.20 $\pm$0.09 12.63 $\pm$0.05 52.53 $\pm$0.21 34.61 $\pm$0.43 23.82 $\pm$0.51 56.67 $\pm$0.21
Dinov2 T5 32.20 $\pm$0.12 16.19 $\pm$0.2 69.98 $\pm$0.32 22.98 $\pm$0.01 11.12 $\pm$0.04 48.28 $\pm$0.11 45.61 $\pm$0.22 34.01 $\pm$0.27 70.01 $\pm$0.11
BUDDY CLIP 31.55 $\pm$0.13 15.05 $\pm$0.43 70.92 $\pm$0.25 23.44 $\pm$0.26 11.06 $\pm$0.20 51.08 $\pm$0.5 43.25 $\pm$0.23 31.84 $\pm$0.35 67.93 $\pm$0.03
ViT T5 30.41 $\pm$0.40 14.11 $\pm$0.28 69.55 $\pm$0.80 22.82 $\pm$0.19 10.24 $\pm$0.12 51.04 $\pm$0.39 43.18 $\pm$0.53 31.73 $\pm$0.54 67.89 $\pm$0.57
ImageBind 33.02 $\pm$0.44 17.61 $\pm$0.43 69.17 $\pm$0.43 24.35 $\pm$0.24 12.05 $\pm$0.46 51.44 $\pm$0.87 41.56 $\pm$0.61 29.89 $\pm$0.91 67.41 $\pm$0.05
Dinov2 T5 30.02 $\pm$0.34 13.78 $\pm$0.19 69.18 $\pm$0.67 22.95 $\pm$0.06 10.45 $\pm$0.09 50.87 $\pm$0.61 43.25 $\pm$0.13 31.77 $\pm$0.33 68.08 $\pm$0.42
MLP CLIP 28.22 $\pm$0.09 14.54 $\pm$0.16 59.40 $\pm$0.08 21.10 $\pm$0.04 10.70 $\pm$0.03 42.77 $\pm$0.05 11.03 $\pm$0.06 4.87 $\pm$0.04 21.61 $\pm$0.11
ViT T5 24.81 $\pm$0.05 11.63 $\pm$0.05 54.78 $\pm$0.04 17.65 $\pm$0.06 8.14 $\pm$0.04 36.77 $\pm$0.06 11.10 $\pm$0.17 4.84 $\pm$0.15 21.94 $\pm$0.24
ImageBind 30.45 $\pm$0.14 15.91 $\pm$0.10 64.10 $\pm$0.07 22.18 $\pm$0.02 11.42 $\pm$0.04 44.86 $\pm$0.06 7.73 $\pm$0.06 3.37 $\pm$0.07 13.26 $\pm$0.03
Dinov2 T5 24.81 $\pm$0.16 11.62 $\pm$0.18 54.97 $\pm$0.22 17.53 $\pm$0.11 8.07 $\pm$0.09 36.53 $\pm$0.26 10.28 $\pm$0.04 4.49 $\pm$0.05 19.86 $\pm$0.03
4.3 Dataloader and Evaluators
MM-Graph provides a standardized dataloader and evaluators implementation for all five datasets. The dataloader is built upon PyTorch’s DataLoader class and the Deep Graph Library (DGL) . It efficiently handles the creation and transmission of graph samples, allowing for mini-batch training of GNNs. The evaluator is based on OGB’s evaluator and performs standardized and reliable evaluation.
The dataloader supports various sampling strategies, such as neighbor sampling and layer-wise sampling, to scale the training process to large graphs. It also enables distributed training using PyTorch’s distributed data parallel (DDP) for faster training on multiple GPUs.
By standardizing the GNNs, feature encoders, dataloader and evaluators, MM-Graph ensures a fair and comprehensive evaluation of graph learning algorithms on multimodal graph data.
5 Empirical Analysis
We conduct a comprehensive evaluation of our proposed benchmark on both link prediction and node classification tasks. Our experiments aim to answer the following research questions:
(i) Are multimodal-GNNs better than conventional GNNs on MM-Graph?
(ii) What is the best approach for encoding features? Is alignment between modalities necessary?
(iii) How much performance gain can be achieved by using multimodal features compared to text features alone?
To ensure fair and rigorous comparisons between various feature encoders and GNNs, we adopt consistent experimental settings across all experiments. We perform automatic hyperparameter tuning using Optuna [1] to optimize the performance of each model. Detailed experimental setup and hyperparameters are provided in the Appendix C.
Figure 2: Feature alignment is important. We compare the performance of various feature encoders, find aligned features, e.g. CLIP and Imagebind result in much better performance compared with unaligned features on Amazon-Sports and Amazon-Cloth. Among them, Imagebind performs the best across backbones. This indicates the importance of using aligned features on these datasets.
5.1 Results
The detailed link prediction results on Amazon-Sports, Amazon-Cloth and Goodreads-LP is shown in Table  3. The detailed node classification results on is shown in Table  4. Here are our findings:
Multimodal GNNs perform poorly on the proposed MM-Graph datasets. As shown in Fig  5.1, we observe that MMGCN and MGAT, the two multimodal GNNs specifically designed to process multimodal input data does not perform better than conventional GNN models such as SAGE. As MMGCN and MGAT mostly just perform message passing and aggregation on each modality separately and then fuse the information across modality together at last, we hypothesize that it is likely to do the insufficient fusion of modalities, and calls for better model designs of multimodal structured data.
Feature Encoders (Alignment). CLIP and ImageBind are two feature encoders which can map features of various modalities to the same embedding space while T5ViT and T5Dinov2 uses SOTA embedding methods for each modality separately, without any specific alignment layers. Interesting, we found the results to vary across datasets. As shown in Figure  2, we found that for Amazon-Sports and Amazon-Cloth, the two amazon datasets, aligned multimodal features performs consistently better than unaligned features. Specifically, we also found that ImageBind consistently gives the best results across backbones. On Goodreads-LP, we found that T5Dinov2 performs best most of the times, this indicates that Dinov2 excels other vision models in the Optical Character Recoginition (OCR) task.
Table 4: Node classification results on Goodreads-NC and Ele-Fashion. *OOM = out of GPU memory.
Image Encoder Text Encoder Ele-fashion Goodreads-NC
MMGCN CLIP CLIP 86.10 $\pm$0.50 OOM
T5 ViT 82.39 $\pm$0.30 OOM
ImageBind ImageBind 86.21 $\pm$0.94 OOM
T5 Dinov2 85.53 $\pm$0.33 OOM
MGAT CLIP CLIP 84.66 $\pm$0.29 OOM
T5 ViT 84.01 $\pm$0.08 OOM
ImageBind ImageBind 86.12 $\pm$0.08 OOM
T5 Dinov2 84.54 $\pm$0.27 OOM
GCN CLIP CLIP 79.83 $\pm$0.03 81.61 $\pm$0.01
T5 ViT 79.63 $\pm$0.07 81.67 $\pm$0.03
ImageBind ImageBind 80.35 $\pm$0.02 78.91 $\pm$0.04
T5 Dinov2 79.37 $\pm$0.04 81.71 $\pm$0.03
SAGE CLIP CLIP 87.10 $\pm$0.02 83.30 $\pm$0.02
T5 ViT 84.41 $\pm$0.09 83.03 $\pm$0.04
ImageBind ImageBind 87.71 $\pm$0.13 80.39 $\pm$0.21
T5 Dinov2 85.31 $\pm$0.09 82.99 $\pm$0.08
MLP CLIP CLIP 85.16 $\pm$0.03 72.29 $\pm$0.02
T5 ViT 84.98 $\pm$0.05 67.82 $\pm$0.07
ImageBind ImageBind 88.73 $\pm$0.01 58.75 $\pm$0.05
T5 Dinov2 84.87 $\pm$0.01 68.83 $\pm$0.03
Figure 3: Multimodal GNNs underperforms conventional GNNs. We compare the best performance of multimodal GNNs (MMGCN/MGAT) and conventional GNNs (SAGE, GCN, BUDDY). Conventional GNNs consistently perform better across datasets, which justifies the importance of building MM-Graph  and calls for better multimodal GNN designs.
[…]
5.2 Challenges and Future Directions.
First, existing multimodal GNNs are all based on full-batch training schemes and it would go out of GPU memory when the graph is large, e.g. Goodreads-NC and Goodreads-LP. Mini-batch multimodal GNNs are worth exploring in the future. Second, we find that it is hard to find a graph dataset with audio/video features with open-source access, even though music/video recommendation is an task of great importance in real-world deployed systems. We believe the study is similar to what is conducted here and the ImageBind features we use can be easily extended to these scenarios. However, integrating multimodal graphs with audio/video features still require future exploration.
Figure 4: Multimodal features are helpful for graph learning. Multimodal features, performs better than text features across datasets and tasks, which justifies the necessity of introducing multimodal graph datasets.
6 Conclusion
In summary, this paper introduces the Multimodal Graph Benchmark (MM-Graph), a comprehensive benchmark for multimodal graph data. MM-Graph consists of five diverse datasets spanning various domains and tasks, featuring rich textual and visual node features. To enable a fair and extensive evaluation, MM-Graph standardizes the GNN architectures, feature encoders, dataloaders and evaluators used across all datasets. The GNNs include both unimodal and multimodal architectures, while the feature encoders leverage state-of-the-art models for capturing semantic information from text and images. Experiments on MM-Graph datasets provide valuable insights into the performance of different GNN and feature encoder combinations in multimodal graph learning settings. By offering a unified evaluation framework, MM-Graph aims to facilitate research on multimodal graph representation learning and drive the development of more powerful and robust graph learning algorithms for real-world applications.
Social impacts. MM-Graph has the potential to significantly advance research on multimodal graph representation learning, which can enable novel applications that positively transform various aspects of society. Many real-world domains, such as healthcare, social networks, and urban computing, involve multimodal graph data. By providing a comprehensive evaluation framework, MM-Graph can accelerate the development of more powerful graph learning algorithms for these domains. MM-Graph also promotes reproducible research by offering standardized benchmarks, fostering collaboration within the research community. Ultimately, MM-Graph aims to unlock the value of multimodal graph data to enable impactful real-world applications. However, as with all technologies of this nature, it is important to consider how it can be used responsibly and with responsible development in mind. There are risks of these algorithms perpetuating or amplifying societal biases if the underlying graph data contains biased patterns. Additionally, misuse of multimodal graph can cause disinformation and targeted manipulation. It is therefore crucial that researchers using MM-GRAPH prioritize the responsible development of multimodal graph algorithms in line with ethical principles.
Appendix A Graph Statistics Metrics
To gain insights into the structural properties of the graphs in each dataset, we compute the following graph statistics metrics:
  • Average Degree. This metric captures the density of the graphs, with higher values indicating denser graphs. It is calculated as the average number of edges connected to each node in the graph.
  • Average Clustering Coefficient (CC). The clustering coefficient measures the tendency of nodes to form tightly connected groups or clusters. A higher clustering coefficient suggests a greater likelihood of nodes forming densely connected communities. It is computed as the average of the local clustering coefficients of all nodes in the graph.
  • Resource Allocation. Resource Allocation is a commonly used heuristic for link prediction tasks. It measures the likelihood of two nodes being connected based on their shared neighbors. The performance of this heuristic on each dataset serves as an indication of how well simple heuristic-based methods may perform, providing a baseline for more sophisticated graph learning algorithms.
  • Transitivity. Transitivity quantifies the probability that two nodes with a common neighbor are also connected to each other, forming a triangle. It is calculated as the ratio of the number of closed triplets (triangles) to the total number of triplets (both open and closed) in the graph.
  • Edge Homophily. Edge Homophily captures the similarity of node labels with respect to the labels of their neighboring nodes. A high Edge Homophily suggests that nodes with similar labels tend to connect to each other, which can be exploited by graph learning algorithms. Edge Homophily is not applicable (N/A) to link prediction datasets, as nodes in these datasets do not have labels.
Appendix B Link Prediction Evaluation
Negative Sampling. For the task of link prediction, we need to evaluate the performance of models in ranking positive edges (existing edges in the graph) higher than negative edges (non-existent edges). To generate negative edges for evaluation, we adopt the widely-used approach of dedicatedly sampling a fixed number of negative edges (e.g., 150) for each positive edge. This approach has been shown to be more effective than using a single, large set of negative edges for all positive edges [18].
Given a set of positive edges and their corresponding negative edges, we employ the following evaluation metrics to assess the link prediction performance:
Hits@K. This metric measures whether the true positive edge is ranked within the top K predictions made by the link prediction model. Specifically, the model scores and ranks all positive and negative edges. Hits@K is then calculated as the fraction of positive edges that are ranked among the top K predictions.
A higher Hits@K value indicates better performance in ranking positive edges among the top predictions. However, it does not provide information about the specific ranks of positive edges beyond the top K. Therefore, we also report the Mean Reciprocal Rank (MRR) metric.
Mean Reciprocal Rank (MRR). MRR is a widely-used metric that considers the specific ranks of positive edges in the predicted ranking. For each positive edge, the reciprocal of its rank in the predicted ranking is calculated. MRR is then computed as the mean of these reciprocal ranks over all positive edges.
A higher MRR value indicates better ranking performance, with a perfect ranking yielding an MRR of 1. MRR provides a more nuanced evaluation of the ranking quality compared to Hits@K, as it considers the entire ranking rather than just the top K predictions.
By reporting both Hits@K and MRR, we provide a comprehensive evaluation of link prediction performance on the multimodal graph datasets in MM-Graph. These metrics enable researchers to assess the strengths and weaknesses of different graph learning algorithms in ranking positive edges higher than negative edges, a crucial task in many real-world applications involving graph data.
Appendix C Experimental details
To ensure a fair and comprehensive evaluation of different graph learning algorithms on the Multimodal Graph Benchmark (MM-Graph), we conduct extensive experiments with rigorous experimental settings and hyperparameter tuning. In this section, we provide detailed information about the experimental setup, hyperparameter search process, and optimization strategies employed in our study.
Table 5: Best performing hyperparameters under Optuna search on Amazon-Sports, Amazon-Cloth and Goodreads-LP. *OOM = out of GPU memory.
Encoder Amazon-Sports Amazon-Cloth Goodreads-LP
Image Text LR # Layers LR # Layers LR # Layers
MMGCN CLIP 0.002595055216 1 0.00120638677883 2 OOM OOM
ViT T5 0.00372938084 1 0.00153969042325 2 OOM OOM
ImageBind 0.0008025598304 1 0.00051054426806 1 OOM OOM
Dinov2 T5 0.001184411766 2 0.00157538063161 2 OOM OOM
MGAT CLIP 0.001706100258 1 0.000109845035707 3 OOM OOM
ViT T5 0.001673818621 3 0.000242709240618 3 OOM OOM
ImageBind 0.000221326243 2 0.0001492753647 2 OOM OOM
Dinov2 T5 0.0001216520274 2 0.00007883247879 3 OOM OOM
GCN CLIP 0.00199258857 1 0.001250668422 1 0.0004599971993 1
ViT T5 0.0004748389636 1 0.0003867480145 1 0.0002695086535 1
ImageBind 0.0002173991668 1 0.00025841621630 1 0.0003440800285 1
Dinov2 T5 0.0003867480145 1 0.00050429857558 1 0.0005790300826 2
SAGE CLIP 0.000888801759 3 0.0006242779054 3 0.001115782634 3
ViT T5 0.0009069080338 3 0.0009159332036 3 0.001115782634 3
ImageBind 0.0007536062461 3 0.001115782634 3 0.0004026026641 3
Dinov2 T5 0.0009159332036 3 0.0009159332036 3 0.001115782634 3
BUDDY CLIP 0.0001996185501 1 0.0001073125661 2 0.0003131657128 1
ViT T5 0.0001782348253 1 0.000104734288 1 0.000101783244 1
ImageBind 0.0001082982111 3 0.0001048823969 3 0.001856957299 1
Dinov2 T5 0.0002695086535 1 0.0001385185507 1 0.0001996185501 1
MLP CLIP 0.0005982646116 1 0.0004259767004 1 0.0002317306399 3
ViT T5 0.00027681242 3 0.0003867480145 1 0.0001156508232 3
ImageBind 0.0001511847491 2 0.0001730467441 2 0.000101783244 3
Dinov2 T5 0.0002317306399 3 0.0003867480145 1 0.0001156508232 3
Table 6: Best performing hyperparameters under Optuna search on Ele-fashion and Goodreads-NC. *OOM = out of GPU memory.
Encoder Ele-fashion Goodreads-NC
Image Text LR # Layers LR # Layers
MMGCN CLIP 0.01617331603 1 OOM OOM
ViT T5 0.0001215688562 2 OOM OOM
ImageBind 0.01590607512 2 OOM OOM
Dinov2 T5 0.01233773648 1 OOM OOM
MGAT CLIP 0.0003211703387 1 OOM OOM
ViT T5 0.0004549756184 2 OOM OOM
ImageBind 0.003739550104 2 OOM OOM
Dinov2 T5 0.0002344834419 1 OOM OOM
GCN CLIP 0.002747501509 2 0.0004599971993 1
ViT T5 0.0007123650663 2 0.0002695086535 1
ImageBind 0.0007123650663 2 0.0003440800285 1
Dinov2 T5 0.0007123650663 2 0.0005790300826 2
SAGE CLIP 0.003235718844 3 0.001115782634 3
ViT T5 0.001523236953 2 0.001115782634 3
ImageBind 0.00174700527 2 0.0004026026641 3
Dinov2 T5 0.00134217448 2 0.001115782634 3
MLP CLIP 0.004407317752 3 0.00174700527 3
ViT T5 0.001342211088 2 0.0004414532347 3
ImageBind 0.00278539483 2 0.0001409760217 3
Dinov2 T5 0.001287758261 3 0.0004611586099 3
C.1 Hyperparameter Tuning
Proper hyperparameter tuning is crucial for obtaining reliable and meaningful results when evaluating machine learning models. To this end, we perform automatic hyperparameter search using Optuna [1], a state-of-the-art hyperparameter optimization framework. The hyperparameter search is directly optimized towards maximizing the evaluation metric of interest for each task.
Specifically, for the link prediction task, we optimize the hyperparameters to maximize the Mean Reciprocal Rank (MRR) metric, as it provides a more nuanced evaluation of ranking performance compared to Hits@K. For the node classification task, we optimize the hyperparameters to maximize the Accuracy metric. The hyperparameter search space includes the following key hyperparameters:
  • Learning Rate (LR): We explore a wide range of learning rates, from 1e-1 to 1e-5, to find the optimal value for each model and dataset combination.
  • Number of GNN Layers (# Layers): We vary the number of Graph Neural Network (GNN) layers from 1 to 3, as the depth of the GNN architecture can significantly impact its performance.
To ensure robust and reliable results, we perform 20 independent hyperparameter studies for each combination of feature encoders and GNN models on a single A40 GPU. Each experiment is run three times with different random seeds, and we report the mean and standard deviation of the respective performance metrics.
C.2 Optimization
All experiments are run with 3 different random seeds to account for the stochasticity in model training and initialization. We report the mean and standard deviation of the respective performance metrics (MRR for link prediction and Accuracy for node classification) across the 3 runs.
For optimization, we use the Adam optimizer with default settings, which has been shown to work well for a wide range of deep learning tasks. To further improve the learning process, we employ a learning rate scheduler with a decay factor of $\gamma=0.1$ and a step size of 5 epochs. This allows the model to make larger updates in the early stages of training and fine-tune the parameters in the later stages.
The best-performing hyperparameters for each combination of feature encoders and GNNs are presented in Table 5 and 6. These hyperparameters are selected based on the highest mean performance across the 3 runs.
C.3 Reproducibility
We make the source code for MM-Graph publicly available on GitHub, along with detailed documentation and instructions for running the experiments. The datasets used in MM-Graph are also made accessible through the Hugging Face Datasets platform, allowing researchers to easily access and use the data for their own experiments.
By providing a comprehensive experimental setup, detailed hyperparameter configurations, and open-source code and datasets, we aim to foster reproducibility and encourage further research on multimodal graph learning using the MM-Graph benchmark.
Appendix D Dataset
In this section, we provide details on the release and maintenance of the Multimodal Graph Benchmark (MM-Graph) datasets, the file structure of the datasets, and instructions on how to use them effectively.
D.1 Dataset Release
To facilitate easy access and widespread use of the MM-Graph datasets, we have created a dedicated webpage at https://mm-graph-benchmark.github.io/. The webpage provides a direct link to the datasets and showcases representative examples from each dataset, allowing researchers to quickly grasp the nature and diversity of the data.
The MM-Graph datasets are hosted on the Hugging Face Datasets platform at https://huggingface.co/datasets/mm-graph-org/mm-graph. Hugging Face Datasets is a popular and reliable platform for hosting and sharing datasets in the machine learning community. By leveraging this platform, we ensure the long-term maintenance and preservation of the MM-Graph datasets. We plan to periodically update the repository to address any issues or incorporate feedback from the community.
The MM-Graph datasets are released under the Apache 2.0 license, which allows for open access and permissive use of the data for research purposes. This licensing choice aligns with our goal of fostering collaborative research and encouraging the development of advanced multimodal graph learning algorithms.
To complement the datasets, we also provide a standardized benchmark, evaluator, and various GNN backbones and feature encoders on our GitHub repository at https://github.com/mm-graph-benchmark/mm-graph-benchmark. The repository includes detailed documentation and example usage to facilitate the adoption and extension of the MM-Graph benchmark.
D.2 Dataset File Structure
The MM-Graph directory contains five separate folders, each corresponding to a specific dataset: amazon-sports, amazon-cloth, goodreads-lp, goodreads-nc, and ele-fashion. Each folder contains the necessary files for the respective dataset. For the three link prediction datasets (amazon-sports, amazon-cloth, and goodreads-lp), each folder contains the following files:
  • lp-edge-split.pt: This file contains the graph structures, along with the train, validation, and test edges, as well as their corresponding negative pairs.
  • clip_feat.pt, t5vit_feat.pt, imagebind_feat.pt, t5dino_feat.pt: These files contain the node features obtained from various feature encoders, as discussed in Section 4.2 of the paper.
For the two node classification datasets (goodreads-nc and ele-fashion), each folder contains the following files:
  • edges.pt: This file contains the graph structures.
  • split.pt: This file indicates the train, validation, and test nodes.
  • label-w-missing.pt: This file stores the node labels. Nodes with missing labels are assigned to a dummy class and are not used as train, validation, or test nodes.
  • clip_feat.pt, t5vit_feat.pt, imagebind_feat.pt, t5dino_feat.pt: These files contain the node features obtained from various feature encoders, as discussed in Section 4.2 of the paper.
D.3 Dataset Usage
To load and use the MM-Graph datasets, we provide a user-friendly Python package.
For detailed examples and usage instructions, please refer to the GitHub repository. The repository provides comprehensive documentation, including installation instructions, dataset descriptions, and example scripts for training and evaluating models on the MM-Graph datasets.
By providing a standardized and easy-to-use interface for accessing the datasets, along with detailed documentation and examples, we aim to facilitate the adoption of the MM-Graph benchmark and enable researchers to focus on developing and evaluating their multimodal graph learning models.

### v2

@@
-Multimodal Graph Benchmark
+Mosaic of Modalities: A Comprehensive Benchmark for Multimodal Graph Learning
 Abstract
-Associating unstructured data with structured information is crucial for real-world tasks that require relevance search.
-However, existing graph learning benchmarks often overlook the rich semantic information associate with each node.
-To bridge such gap, we introduce the Multimodal Graph Benchmark (MM-Graph), the first comprehensive multi-modal graph benchmark that incorporates both textual and visual information.
-MM-Graph surpasses previous efforts, which have primarily focused on text-attributed graphs with various connectivity patterns.
-MM-Graph consists of five graph learning datasets of various scales that are appropriate for different learning tasks.
-Their multimodal node features, enabling a more comprehensive evaluation of graph learning algorithms in real-world scenarios.
-To facilitate research on multimodal graph learning, we further provide an extensive study on the performance of various graph neural networks in the presence of features from various modalities.
-MM-Graph aims to foster research on multimodal graph learning and drive the development of more advanced and robust graph learning algorithms.
-By providing a diverse set of datasets and benchmarks, MM-Graph enables researchers to evaluate and compare their models in realistic settings, ultimately leading to improved performance on real-world applications that rely on multimodal graph data.
+Graph machine learning has made significant strides in recent years, yet the integration of visual information with graph structure and its potential for improving performance in downstream tasks remains an underexplored area.
+To address this critical gap, we introduce the Multimodal Graph Benchmark (MM-Graph), a pioneering benchmark that incorporates both visual and textual information into graph learning tasks.
+MM-Graph extends beyond existing text-attributed graph benchmarks, offering a more comprehensive evaluation framework for multimodal graph learning.
+Our benchmark comprises seven diverse datasets of varying scales (ranging from thousands to millions of edges), designed to assess algorithms across different tasks in real-world scenarios.
+These datasets feature rich multimodal node attributes, including visual data, which enables a more holistic evaluation of various graph learning frameworks in complex, multimodal environments.
+To support advancements in this emerging field, we provide an extensive empirical study on various graph learning frameworks when presented with features from multiple modalities, particularly emphasizing the impact of visual information.
+This study offers valuable insights into the challenges and opportunities of integrating visual data into graph learning.
 1 Introduction
-Graphs are ubiquitous data structures that can effectively represent complex relationships and interactions between entities in various domains, such as social networks, biological systems, and recommendation systems.
-In real-world scenarios, these entities often possess rich semantic information in the form of unstructured data, such as images and text descriptions.
-Associating these unstructured data with the structured graph information is essential for tasks that require relevance search and information retrieval, such as recommendation systems [27, 3, 10].
-Despite the importance of multimodal graph learning, existing graph learning benchmarks typically focus on investigating graphs with various connectivity patterns, such as long-range links, out-of-distribution scenarios, etc [5, 9, 19, 20, 13].
-These benchmarks aim to evaluate the performance of graph neural networks (GNNs) and other graph learning algorithms on different graph structures and topologies.
-However, these benchmarks under-utilize the rich semantic information naturally present in each node, limiting their ability to evaluate graph learning algorithms in realistic settings.
-With the prevalence of language modeling, there has been an outbreak in the effective integration of text and graph topology on text-attributed graphs, leading to a comprehensive benchmark on text-attributed graphs [33, 35, 15, 23].
-While textual information undoubtedly plays a crucial role in understanding the data and making predictions for user-centric tasks, it is essential to recognize that other modalities, particularly visual information, can be helpful and provide complementary information that text alone cannot capture.
-For example, as shown in Figure  1(a), visually similar products may have distinct text features and ignoring such semantic gap between modalities may prevent GNNs from fully exploiting the relationships between multimodal semantics and structures.
-Figure 1: Visualization of our Multimodal Graph Benchmark.
-All nodes of our benchmark have both visual and text features.
-(a) Amazon-Sports: The image and text come from the original image and title of the sports equipment.
-(b) Goodreads-LP: The image comes from the cover of the book.
-We do not show the text features of Goodreads-LP since the book description is very long.
-(c) Ele-fashion: The image and text come from the original image and title of the fashion product.
-To address this limitation, we introduce the Multimodal Graph Benchmark (MM-Graph), the first comprehensive graph benchmark that incorporates both textual and visual information.
-MM-Graph consists of five graph learning datasets of various scales, designed for diverse downstream tasks such as node classification and link prediction.
-These datasets contain multimodal node features, including text and images, alongside the traditional graph structure.
-By providing a diverse set of datasets and benchmarks, MM-Graph enables researchers to evaluate and compare their models in realistic settings, fostering research on multimodal graph learning.
-MM-Graph also standardizes GNN architectures, feature encoders, and dataloaders and evaluators in order to enable a comprehensive evaluation on multimodal graph data
-To show the importance of these datasets in for advancing multimodal graph learning , we adapt two state-of-the-art multimodal graph neural networks from the recommendation systems to the graph learning domain.
-These models have demonstrated strong performance in leveraging multimodal information for recommendation tasks.
-By applying them to the datasets in MM-Graph, we aim to provide a comprehensive evaluation of their effectiveness in various graph learning scenarios.
-We conduct an extensive study on the performance of MMGCN [31], MGAT [28], and other popular graph neural network models, such as GraphSAGE [11], in the presence of multimodal node features.
-Surprisingly, we find that GNNs specifically designed for multimodal scenarios, such as MMGCN and MGAT, do not consistently outperform general-purpose GNN models like GraphSAGE across all datasets in MM-Graph.
-This finding highlights the need for a standard multimodal graph benchmark to rigorously evaluate and compare the performance of different models.
-MM-Graph fills this gap by providing a diverse set of datasets with varying scales, tasks, and modality combinations.
-By establishing a common ground for evaluation, MM-Graph enables researchers to gain a deeper understanding of the strengths and limitations of existing multimodal graph learning algorithms and identify areas for improvement.
-Moreover, MM-Graph opens up new research opportunities in multimodal graph learning.
-The availability of datasets with rich textual and visual information allows researchers to explore novel architectures and techniques that can effectively capture and integrate the semantic information from multiple modalities.
-This can lead to the development of more advanced and robust graph learning algorithms that can handle the complexities of real-world multimodal graph data.
+Graphs are powerful data structures that capture complex relationships across various domains, from social networks to recommendation systems.
+In real-world applications, these graph entities often possess rich, multimodal semantic information, including crucial visual data.
+The integration of visual information with structured graph data is particularly vital for tasks such as relevance search and information retrieval in recommendation systems [38, 4, 12].
+Existing graph learning benchmarks have made significant strides in evaluating graph neural networks (GNNs) on various connectivity patterns and topologies [6, 11, 24, 27, 16], they often overlook the critical role of visual information.
+Moreover, recent advancements in language modeling have led to effective integration of textual information with graph structures, resulting in comprehensive benchmarks for text-attributed graphs [47, 50, 18, 30].
+However, visual information offers unique and invaluable insights that text and structure alone cannot capture [44, 45].
+Visual data can reveal similarities and relationships in artist styles that may not be apparent from textual or structural information alone.
+For instance, as illustrated in LABEL:fig:teaser, visually similar products may have distinct textual descriptions, creating a semantic gap that text-only GNNs might fail to bridge.
+By incorporating visual information, graph learning algorithms can better understand and exploit the intricate relationships between multimodal semantics and graph structures.
+This visual context can significantly enhance the performance of GNNs.
+Integrating visual data into graph learning frameworks enables a more holistic understanding of the underlying relationships in complex, real-world networks.
+It allows GNNs to capture nuanced information that may be missed when relying solely on textual or structural data, leading to more accurate and contextually relevant predictions.
+To address the critical need for integrating visual data in graph machine learning, we introduce the Multimodal Graph Benchmark (MM-Graph), the first comprehensive graph benchmark that explicitly incorporates visual information alongside textual data.
+MM-Graph comprises seven diverse graph learning datasets of varying scales, specifically designed to evaluate two fundamental categories of graph tasks: node and link property predictions.
+These tasks require models to make predictions at different levels of graph granularity, leveraging the rich multimodal information available.
+A key feature of MM-Graph is its inclusion of high-resolution visual node features, which sets it apart from existing benchmarks.
+Each dataset within MM-Graph contains not only textual information and traditional graph structures but also associated images for nodes.
+This visual data provides crucial additional context that can significantly enhance the performance of graph learning algorithms in real-world scenarios.
+While multimodal knowledge graphs (MMKGs) incorporating visual information exist, the community faces challenges with limited datasets, often with outdated or inaccessible image URLs and poor dataset quality [26, 36, 25, 3].
+Recognizing the critical need for high-quality MMKG datasets that effectively integrate visual data [25], our MMKGs are sourced from CoDEx, a recent comprehensive knowledge graph benchmark known for its diverse and interpretable content and challenging link prediction tasks [36].
+By enhancing CoDEx with curated visual information, our MMKGs offer higher quality datasets that address a significant gap in the MMKG community, providing a robust foundation for evaluating how graph learning models incorporate and benefit from visual features in knowledge graph scenarios.
+We highlight the key differences between MM-Graph and existing benchmarks in  Tab.
+1.
+Table 1: Qualitative comparison of existing benchmarks. Abbreviations: LP = Link Prediction. NC = Node Classification. KGC = Knowledge Graph Completion.
+Benchmarks Language Vision NC LP KGC
+GL-Bench [23]
+CS-TAG [47]
+VTKG [20]
+CoDEx [35]
+OGB [16]
+MM-Graph
 Contributions and novelty.
 In this work, we introduce a multimodal graph benchmark known as MM-Graph.
-As design principles, we strive to create non-trivial, realistic data that comes from real-world applications and will benefit real-world deployments.
-Specifically, MM-Graph contains three link prediction datasets and two node classification datasets, as shown in Table  1.
-We summarize our novel contributions as follows: (1) To the best of our knowledge, no existing standardized graph benchmark provides both text and image features.
-This is important as images provide complementary context and details.
-While one can extract keywords and intent from text, images supply visual cues about objects, scenes, emotions, etc.
-that give richer understanding, especially for more abstract or complex tasks.
-(2) We provide a comprehensive analysis on the performance of various GNNs on MM-Graph, highlighting the need of developing new multimodal GNNs.
-(3) In terms of graph benchmarks, MM-Graph is the first benchmark to include not only various GNNs but also different ways of encoding node features, which presents an important and practical question of which feature encoder gives the best features for structure learning tasks.
-Reproducibility: MM-Graph code, datasets and details are publicly available. The datasets are hosted on Huggingface. The evaluation code and baselines are hosted on GitHub with documentations.
-Table 1: Overview of the datasets in the proposed MM-Graph.
+MM-Graph fills this gap by providing a diverse set of datasets with varying scales, tasks, and modality combinations.
+By establishing a common ground for evaluation, MM-Graph enables researchers to reflect on the strengths and limitations of existing multimodal graph learning algorithms and identify areas for improvement.
+As design principles, we strive to create non-trivial, realistic data that comes from real-world applications and can benefit real-world deployments.
+Specifically, MM-Graph contains three link prediction datasets, two node classification datasets, and two knowledge graph completion datasets as shown in Tab.
+2.
+We summarize our novel contributions as follows:
+(1) Text- and Image- Attributed Graph(TIAG) Benchmarks.
+First standardized benchmark providing both textual and visual features across various graph learning tasks, enabling a more holistic understanding of complex data relationships.
+(2) Benchmarking Multimodal GNNs. Comprehensive performance analysis of diverse methods on MM-Graph, highlighting the need for more advanced multimodal Graph Neural Networks (GNNs).
+(3) Exploring Multimodal Feature Encodings.
+Pioneering investigation of various node feature encoding strategies (e.g.
+ImageBind), addressing the challenge of optimal feature representation in graph learning tasks.
+Table 2: Overview of the datasets in our proposed MM-Graph benchmark.
 We present five multimodal graph datasets with varying scale and tasks.
-All datasets are splitted in random.
-LP = Link Prediction.
+All datasets are split at random.
+Abbreviations: LP = Link Prediction.
 NC = Node Classification.
-Name Text Features Visual Features Proposed Task Metrics Scale Split Ratio
+KGC = Knowledge Graph Completion.
+OR = Original.
+Name Text Features Visual Features Task Metrics Scale Split Ratio
 Amazon-Sports Product Titles Product Images LP MRR, Hits@K Small 8/1/1
@@
 Goodreads-NC Book Description Book Images NC Accuracy Large 6/1/3
+MM-CoDEx-s Entity Description Entity Images KGC MRR, Hits@K Small OR
+MM-CoDEx-m Entity Description Entity Images KGC MRR, Hits@K Medium OR
 2 Related Work
@@
 While most existing multimodal learning approaches focus on 1-to-1 mappings between modalities, real-world data often exhibits more complex many-to-many relationships that can be represented as graphs.
-Multimodal graph learning (MMGL) aims to address this challenge by leveraging graph structures to learn from multimodal data with complicated relations among multiple multimodal neighbors  [6, 34].
-Recent efforts in MMGL have focused on integrating GNNs with pretrained language models (LMs) [33].
-Our work builds upon the recent advances in multimodal feature learning and MMGL by introducing a comprehensive benchmark that incorporates both structured and unstructured modalities.
-We adapt up-to-date multimodal feature learning frameworks and state-of-the-art multimodal GNNs and provide an extensive study on their performance in the presence of features from various modalities.
+Multimodal graph learning (MMGL) aims to address this challenge by leveraging graph structure to learn from multimodal data with complicated relations among multiple multimodal neighbors  [7, 48, 8, 14].
+Recent efforts in MMGL have focused on integrating GNNs/Knowledge Graph Embeddings (KGEs) with pretrained language models (LMs) [47].
+Our work builds upon the recent advances in multimodal feature learning and MMGL by introducing a standardized benchmark that incorporates both structured and unstructured modalities.
 Benchmarks for graph representation learning.
@@
 However, when it comes to learning on graphs with rich features, these benchmarks exhibit notable deficiencies.
-Firstly, a majority of these datasets suffer from the absence of raw feature information, limiting the investigation of attribute modeling’s effectiveness.
-Secondly, these datasets often neglect to explore the impact of feature modeling on GNNs.
-Thus, there is a compelling necessity to construct a comprehensive graph benchmark with rich multimodal features.
+Firstly, these datasets suffer from the absence of raw features, limiting the investigation of attribute modeling.
+Secondly, these datasets often neglect to explore the impact of feature modeling.
+Thus, there is a compelling necessity to construct a comprehensive graph benchmark with rich natural features.
 Yan et al.
-proposed the first standardized graph benchmark with rich textual information, fostering the development of encompassing GNNs with prevalent language models [33].
+proposed a graph benchmark with rich textual information, fostering the development of integrating GNNs with prevalent language models [47].
 However, we argue that it is critical to investigate modalities other than text, especially images in structure learning tasks and proposing the first standarized multimodal graph benchmark to foster the development of research in this direction.
-3 MM-Graph: The Multimodal Graph Benchmark
-In this section, we discuss the proposed 5 real-world datasets in MM-Graph.
-The basic information is shown in Table  1.
-We first provide diverse graph-related statistics so as to give a comprehensive understanding of the structural properties of the datasets.
-This information can guide the selection of appropriate graph learning algorithms and help interpret the performance of different models on each dataset.
+The key differences between MM-Graph and existing benchmarks are highlighted in  Tab.
+1.
+3 MM-Graph: Multimodal Graph Benchmark
+In this section, we present the proposed 7 real-world datasets in MM-Graph.
+The general information about them is given in  Tab.
+2.
+We first introduce the data curation process of all datasets, as well as their license.
+For all datasets described below, we use Beautiful Soup to crawl images associated with the nodes in the graphs [33].
 3.1 Data Curation
-We first introduce the data curation process of all datasets, as long as their license.
-Amazon-Sports Amazon-Sports is a link prediction dataset, based on the Amazon-Review dataset [21, 12], where each node corresponds to a product on Amazon in the sports category and the link captures whether two products are co-purchased together.
+Amazon-Sports/Amazon-Cloth are link prediction datasets, based on the Amazon-Review dataset [28, 15], where each node corresponds to a product on Amazon in the sports category and the link captures whether two products are co-purchased together.
 The text features are the titles of the products and the visual features are the high-resolution raw images of the products.
-In Figure  1(a), we show an example of nodes and edges in Amazon-Sports.
-The visual features are images of tents of various shapes under diverse backgrounds and text features are the title of the tents.
-Specifically, we use the data pre-processing script provided by Patton [16].
-Amazon-Cloth Similar to Amazon-Sports, Amazon-Cloth is constructed according to the co-purchasing information of Amazon products in the clothes category.
-The text features are titles of products, e.g.
-Nike mens Revolution 6 Road Running and the visual features are corresponding images.
-We follow the same data curation procedure.
-Goodreads-LP is based on the Goodreads Book Graph dataset  [29, 30].
-Here we construct the graph in the way where each node corresponds to a book on Goodreads and the link captures if a user who likes this book will like the other book, following  [35].
-The text features are the descriptions of the books and the visual features are the cover of the book.
+In LABEL:fig:teaser(a), we show an example of nodes and edges in Amazon-Sports.
+The visual features are images of tents of various shapes under diverse backgrounds and text features are the titles of the tents.
+We preprocess the data by extracting the meta information of each item as well as their co-purchasing information.
+Goodreads-LP/Goodreads-NC are based on the Goodreads Book Graph dataset  [41, 42].
+Here we construct the graph such that each node corresponds to a book on Goodreads and the link captures if a user who likes this book will like the other book, following  [50].
+The text features are the descriptions of the books and the visual features are the the book covers.
 Nodes without images are removed.
-Goodreads-NC is a node classification dataset, based on the Goodreads dataset [29, 30], where each node corresponds to a book on Goodreads and the link captures if a user who likes this book will like the other book.
-The text features are the descriptions of the books and the visual features are the cover image of the book.
-Nodes without images are removed.
-Ele-fashion Ele-fashion is a node classification dataset, based on the Amazon-Fashion dataset [21, 12], where each node corresponds to a product and the link captures if a user who buy this product will also buy the other product.
+Ele-fashion is a node classification dataset, based on the Amazon-Fashion dataset [28, 15], where each node corresponds to a product and a link captures that the two products are copurchased.
 The text features are the titles of the products and the visual features are the images of the products.
+MM-CoDEx-s/MM-CoDEx-m is a knowledge graph completion dataset, based on the CoDEx-s dataset [36].
+The text features come from the description of the Wikipedia article.
+For visual features, we crawl the corresponding images of the entity (e.g., person/place/item) from Wikipedia.
 Data Availability and Ethics.
 Our benchmark is organized from existing open source data, with proper open source licenses.
-Amazon-Sports and Amazon-Cloth are available with Apache License  [^1] https://github.com/PeterGriffinJin/Patton.
-Goodreads-LP, Goodreads-NC and Ele-fashion are released under MIT License  [^2] https://mengtingwan.github.io/data/goodreads.html.
+Amazon-Sports and Amazon-Cloth and Ele-Fashion are available with Apache License  [^1] https://github.com/PeterGriffinJin/Patton.
+Goodreads-LP, Goodreads-NC, MM-CODEX-S and MM-CODEX-M are released under MIT License  [^2] https://mengtingwan.github.io/data/goodreads.html, https://github.com/tsafavi/codex/tree/master.
 These do not involve interaction with humans or private data.
-Table 2: Detailed graph-based statistics of each dataset proposed in MM-Graph.
+Table 3: Detailed graph-based statistics of datasets proposed in MM-Graph.
 CC = Cluster Coefficient.
 RA = Resource Allocation.
-N/A = Not Applicable (Nodes do not have node labels).
-Detailed definition for each statistics provided in the Appendix.
+N/A = Not Applicable (Nodes do not have class labels).
+Detailed definition for each graph statistics or property are provided in the Appendix.
 Name Nodes Edges Average Degree Average CC Average RA Transitivity Edge Homophily
@@
 Ele-fashion 97,766 199,602 4.08 0.1730 0.1467 0.0560 0.7675
+Table 4: Detailed MMKG statistics of datasets proposed in MM-Graph. (+): Positive (true) triples. (-): Verified negative (false) triples.
+Name Entities Relations Train (+) Valid (+) Test (+) Valid (-) Test (-)
+MM-CoDEx-s 1,383 39 14,298 784 802 1028 1074
+MM-CoDEx-m 7,697 51 47,617 2,628 2,595 4,721 4,746
 3.2 Tasks
@@
 Specifically, for Amazon-Sports and Amazon-Cloth, we generate hard valid/test negatives using HeaRT [22].
-HeaRT is recognized as a better way of generating negatives for link prediction tasks in the community.
+HeaRT is recognized as a better way of generating negatives for link prediction in the community.
 For Goodreads-LP, since it takes more than 120 hours to generate hard negatives using HeaRT, we perform random sampling.
 Each positive edge in the validation/test set is ranked against 150 hard negative edges.
-For evaluation metrics, we report MRR, Hits@10, and Hits@1 - the three most commonly-used evaluation metrics for link prediction [13, 18].
+For evaluation metrics, we report MRR, Hits@10, and Hits@1, the three most commonly-used evaluation metrics for link prediction [16, 22].
 For Amazon-Sports and Amazon-Cloth, edges are randomly split into train/valid/test splits according to 8/1/1 ratio.
@@
 Node Classification.
-The task is to predict the book category from $N$ available categories.
-We report accuracy following  [13, 33].
-For Goodreads-NC, there are 10 categories such as History, Children and comics.
+For Goodreads-NC, the task is to predict the category of items from 10 available categories, such as History, Children and comics.
 For Ele-Fashion, there are 12 categories such as shoes, jewelry and dresses.
-For evaluation metrics, we report accuracy, as it’s the most common metric for node classification[13].
+For evaluation metrics, we report accuracy, as it’s the most common metric for node classification[16, 47].
 Nodes are randomly split into train/valid/test following 6/1/3 ratio.
+Knowledge Graph Completion.
+The task is to predict missing links in a knowledge graph given entities and relations.
+Similar to link prediction, the evaluation is based on how well a model ranks positive test edges over negative test edges.
+Specifically, we use the splits and negatives designed in  [36] to ensure the scope and level of difficulty of the task.
+Entities without multimodal information (i.e., entities that do not have descriptions or images) are filtered out.
+For evaluation metrics, we report MRR, Hits@10, Hits@3 and Hits@1.
 3.3 Statistics
-We provide graph-related statistics are presented in Table  2.
-The detailed descriptions for each metric is provided in the Appendix.
-These statistics provide valuable insights into the structural properties of the graphs in each dataset.
-Amazon-Sports.There are 50,250 nodes with a total of 356,202 edges in Amazon-Sports, which is a small-scale link prediction dataset.
-The small scale makes Amazon-Sports great for trying it at first, as it is computation and memory efficient for most applications.
-The average degree is 14, which indicates that each node have 14 neighbors on average.
-Amazon-Cloth.There are 125,839 nodes with a total of 951,271 edges in Amazon-Cloth.
-Amazon-Cloth constitutes a medium-scale link prediction dataset.
-Since the scale of this dataset is much larger than Amazon-Sports, Amazon-Cloth serves as a solid benchmark for real-world e-commerce based applications.
-The average degree is 15, which similar to Amazon-Sports.
-But the average clustering coefficient is much smaller.
-This indicates that nodes in Amazon-Cloth are more likely to scatter around, instead of forming communities despite its high average degree.
-Goodreads-LP.There are 636,502 nodes with a total of 3,437,017 edges, which constitutes a large-scale link prediction dataset. The average degree is 10.
-Goodreads-NC. There are 685,294 nodes with a total of 7,235,084 edges, which constitutes a large-scale node classification dataset. The average degree is 21, highest among all datasets.
-Ele-fashion.
-There are 97,766 nodes with a total of 199,602 edges, which constitutes a medium-scale node classification dataset.
-The average degree is 4, which is the most sparse dataset among MM-Graph.
+We present detailed statistics for each dataset in Tab.
+3 and Tab.
+4, and additional descriptions of each metric and property are provided in the Appendix.
+These statistics provide valuable insights into the structural properties of the graphs.
+Our multimodal graph benchmark encompasses datasets spanning various scales and characteristics.
+The datasets range from small-scale (MM-CoDEx-s) to large-scale (Goodreads-NC) with varying entity and relation counts.
+The diversity in graph sizes, densities, and average degrees provides a comprehensive testbed for evaluating multimodal graph learning approaches across different computational complexities.
 4 Evaluation Framework
-To enable a comprehensive evaluation on multimodal graph data, MM-Graph standardizes the GNN architectures, feature encoders, and dataloaders and evaluators used across all datasets.
+To enable a comprehensive evaluation on multimodal graph data, MM-Graph standardizes the GNN architectures, KGEs, feature encoders, dataloaders and evaluators used.
 4.1 Graph Neural Networks
@@
 Conventional GNNs.
-We standardizes GCN, SAGE, MLP for both link prediction and node classification.
-Since BUDDY is specifically designed for link prediction, so we do not report its performance on node classification.
+We standardize GCN, SAGE, MLP for both link prediction and node classification.
+Since BUDDY is specifically designed for link prediction, we do not report its performance on node classification.
+For all unimodal GNNs, text and image embeddings are first concatenated together and then passed through GNNs.
 Multimodal GNNs.
 While GNNs have shown strong performance on tasks involving graphs with unimodal features, there has been limited work on developing GNNs that can effectively handle multimodal graph data.
-In order to evaluate the effectiveness of GNNs on multimodal graphs, we adapt two recent architectures from the recommendation systems domain to our standardized multimodal graph benchmarks:
-MMGCN is a GCN-based framework that leverages information interchange between users and micro-videos in multiple modalities to refine their modality-specific representations and model users’ fine-grained preferences.
-It constructs separate user-item graphs for each modality.
+To evaluate the effectiveness of GNNs on multimodal graphs, we adapt two recent architectures from the recommendation systems domain to our standardized multimodal graph benchmarks:
+(1) MMGCN constructs separate user-item graphs for each modality.
 The information interchange of users and items in each modality is encoded using GNNs.
-And modality representations are fused in the prediction layer to predict possible recommendations [31].
-For LP, we use the most common dot product decoder to decode possible links.
-For NC, we stack a 3-layer MLP as transform representations to a number of dimensions equal to the number of node classes.
-MGAT extends the popular GAT architecture to multimodal settings.
-It learns modality-specific node representations which are then combined using a cross-modal attention layer  [28].
+Modality representations are fused in the prediction layer to predict possible recommendations [43].
+For link prediction (LP), we use the most common dot product decoder to decode possible links.
+For node classification (NC), we stack a 3-layer MLP to transform representations to a number of dimensions equal to the number of node classes.
+(2) MGAT extends the popular GAT architecture and learns modality-specific node representations which are then combined using a cross-modal attention layer [39].
 This allows MGAT to weigh different modalities based on their importance.
 Similar to MMGCN, we decode links using dot product for LP and perform NC through a 3-layer MLP.
+4.2 Knowledge Graph Embeddings
+MM-Graph includes two state-of-the-art multimodal Knowledge Graph Embeddings (KGEs): MoSE [49] and VISTA [20].
+MoSE introduces modality-split relation embeddings for each modality, moving away from a single modality-shared embedding, which effectively mitigates modality interference.
+VISTA innovatively incorporates visual and textual representations of entities and relations through sophisticated transformers, utilizing entity encoding, relation encoding, and triplet decoding.
+Since VISTA’s architecture scores entities rather than triples, we report performance using the standard 1vsAll setting instead of hard negatives [20, 34].
+Note that we do not report Graph Neural Network (GNN) performance for Knowledge Graph Completion (KGC), as prior research has demonstrated that GNN architectures can inadvertently conflate scoring functions in Knowledge Graph Embedding, potentially degrading overall performance [21].
 4.3 Feature Encoders
 Text encoders.
-For text encoders, we select CLIP  [24], T5  [25] and ImageBind  [8].
-T5 is selected as the STOA text embedding models, while CLIP and ImageBind are selected so that the output text representations aligns with visual representations.
-Visual encoders.
-For visual encoders, we select CLIP  [24], ViT  [4], ImageBind  [8] and Dinov2  [22].
-ViT and Dinov2 are selected as two different types of visual encoders, while ViT are explicitly trained with supervision, Dinov2 learns robust visual features without supervision.
-CLIP and ImageBind are selected so that the output text representations aligns with visual representations.
-Specifically, Imagebind has potentials to extend multimodal graph learning to other modalities, such as audio and video thanks to its ability to bind all modalities into one embedding space.
-Studies on various feature encoders give insights on design choices of multimodal graph learning, such as (1) how important it is to align multiple modalities into a unified embedding space, (2) if supervised feature encoders work better than unsupervised feature encoders.
-Note that feature encoders are only used to extract features for GNNs and are never trained upon.
-Table 3: Link prediction results on Amazon-Sports, Amazon-Cloth and Goodreads-LP. *OOM = out of GPU memory. Highlighted box indicates the best performing combination for each dataset.
+For text encoders, we select CLIP [31], T5 [32], and ImageBind [10].
+T5 is chosen as the state-of-the-art (SOTA) text embedding model, while CLIP and ImageBind are selected to ensure that the output text representations align with visual representations.
+Specifically, CLIP provides a multi-modal embedding approach that jointly trains text and image encoders to maximize cosine similarity between text and image embeddings.
+ImageBind extends this concept by enabling cross-modal representation learning across multiple modalities.
+Table 5: Link prediction results on Amazon-Sports, Amazon-Cloth and Goodreads-LP.
+Highlighted box indicates the best performing combination for each dataset.
+Conventional GNNs such as SAGE performs best across datasets.
+Aligned feature embeddings, e.g.
+CLIP and ImageBind outperforms unaligned features.
 Encoder Amazon-Sports Amazon-Cloth Goodreads-LP
 Image Text MRR $\uparrow$ H@1 $\uparrow$ H@10 $\uparrow$ MRR $\uparrow$ H@1 $\uparrow$ H@10 $\uparrow$ MRR $\uparrow$ H@1 $\uparrow$ H@10 $\uparrow$
-MMGCN CLIP 31.96 $\pm$0.10 16.35 $\pm$0.11 68.46 $\pm$0.08 22.20 $\pm$0.05 10.76 $\pm$0.1 46.62 $\pm$0.12 OOM OOM OOM
-ViT T5 30.33 $\pm$0.03 15.01 $\pm$0.05 66.41 $\pm$0.11 19.45 $\pm$0.34 9.22 $\pm$0.20 40.49 $\pm$0.61 OOM OOM OOM
-ImageBind 31.74 $\pm$0.21 16.45 $\pm$0.13 67.39 $\pm$0.74 24.72 $\pm$0.19 12.47 $\pm$0.09 51.32 $\pm$0.56 OOM OOM OOM
-Dinov2 T5 30.04 $\pm$0.27 14.98 $\pm$0.07 64.56 $\pm$0.56 21.77 $\pm$0.23 10.47 $\pm$0.12 45.81 $\pm$0.52 OOM OOM OOM
-MGAT CLIP 27.56 $\pm$0.30 13.55 $\pm$0.29 60.21 $\pm$0.21 21.38 $\pm$0.23 10.39 $\pm$0.22 44.60 $\pm$0.36 OOM OOM OOM
-ViT T5 30.15 $\pm$0.34 15.28 $\pm$0.34 64.84 $\pm$0.41 20.59 $\pm$0.41 9.79 $\pm$0.30 43.44 $\pm$0.76 OOM OOM OOM
-ImageBind 30.15 $\pm$0.12 15.50 $\pm$0.05 64.20 $\pm$0.43 22.13 $\pm$0.27 10.96 $\pm$0.15 45.84 $\pm$0.57 OOM OOM OOM
-Dinov2 T5 28.91 $\pm$0.09 14.47 $\pm$0.18 62.11 $\pm$0.22 21.42 $\pm$0.13 10.38 $\pm$0.13 44.11 $\pm$0.50 OOM OOM OOM
+MMGCN CLIP 31.96 $\pm$0.10 16.35 $\pm$0.11 68.46 $\pm$0.08 22.20 $\pm$0.05 10.76 $\pm$0.1 46.62 $\pm$0.12 31.84 $\pm$0.09 18.63 $\pm$0.31 59.85 $\pm$0.19
+ViT T5 30.33 $\pm$0.03 15.01 $\pm$0.05 66.41 $\pm$0.11 19.45 $\pm$0.34 9.22 $\pm$0.20 40.49 $\pm$0.61 31.11 $\pm$0.25 19.30 $\pm$0.45 56.24 $\pm$0.19
+ImageBind 31.74 $\pm$0.21 16.45 $\pm$0.13 67.39 $\pm$0.74 24.72 $\pm$0.19 12.47 $\pm$0.09 51.32 $\pm$0.56 26.32 $\pm$0.23 16.05 $\pm$0.22 46.37 $\pm$0.66
+DINOv2 T5 30.04 $\pm$0.27 14.98 $\pm$0.07 64.56 $\pm$0.56 21.77 $\pm$0.23 10.47 $\pm$0.12 45.81 $\pm$0.52 27.64 $\pm$0.95 16.21 $\pm$0.65 51.46 $\pm$1.71
+MGAT CLIP 27.56 $\pm$0.30 13.55 $\pm$0.29 60.21 $\pm$0.21 21.38 $\pm$0.23 10.39 $\pm$0.22 44.60 $\pm$0.36 74.75 $\pm$1.23 64.53 $\pm$1.48 92.81 $\pm$0.64
+ViT T5 30.15 $\pm$0.34 15.28 $\pm$0.34 64.84 $\pm$0.41 20.59 $\pm$0.41 9.79 $\pm$0.30 43.44 $\pm$0.76 75.26 $\pm$1.21 65.23 $\pm$1.62 92.90 $\pm$1.89
+ImageBind 30.15 $\pm$0.12 15.50 $\pm$0.05 64.20 $\pm$0.43 22.13 $\pm$0.27 10.96 $\pm$0.15 45.84 $\pm$0.57 74.77 $\pm$0.49 64.95 $\pm$0.61 92.51 $\pm$0.47
+DINOv2 T5 28.91 $\pm$0.09 14.47 $\pm$0.18 62.11 $\pm$0.22 21.42 $\pm$0.13 10.38 $\pm$0.13 44.11 $\pm$0.50 74.89 $\pm$1.46 64.70 $\pm$1.98 92.92 $\pm$0.41
 GCN CLIP 31.38 $\pm$0.08 16.58 $\pm$0.13 66.14 $\pm$0.08 22.28 $\pm$0.05 11.83 $\pm$0.04 43.52 $\pm$0.10 25.34 $\pm$0.06 13.81 $\pm$0.12 50.36 $\pm$0.14
@@
 DINOv2 T5 24.81 $\pm$0.16 11.62 $\pm$0.18 54.97 $\pm$0.22 17.53 $\pm$0.11 8.07 $\pm$0.09 36.53 $\pm$0.26 10.28 $\pm$0.04 4.49 $\pm$0.05 19.86 $\pm$0.03
+Visual encoders.
+For visual encoders, we select CLIP [31], ViT [5], ImageBind [10], and DINOv2 [29].
+ViT and DINOv2 represent two distinct approaches to visual feature learning: ViT is explicitly trained with supervision, while DINOv2 learns robust visual features through self-supervised methods.
+CLIP and ImageBind are chosen to explore text-visual representation alignment.
+Notably, ImageBind demonstrates potential for extending multimodal graph learning across diverse modalities like audio and video by embedding them in a unified space.
+Our selection of feature encoders enables us to investigate critical multimodal graph learning design, specifically:
+  • The importance of aligning multiple modalities into a unified embedding space.
+  • Comparative performance of supervised versus unsupervised feature encoders.
 4.4 Dataloader and Evaluators
-MM-Graph provides a standardized dataloader and evaluators implementation for all five datasets.
-The dataloader is built upon PyTorch’s DataLoader class and the Deep Graph Library (DGL) .
-It efficiently handles the creation and transmission of graph samples, allowing for mini-batch training of GNNs.
-The evaluator is based on OGB’s evaluator and performs standardized and reliable evaluation.
+MM-Graph also provides a standardized dataloader and evaluator implementation.
+The dataloader is built upon PyTorch’s DataLoader class.
+The evaluator is based on PyTorch’s evaluator and performs standardized and reliable evaluation.
 The dataloader supports various sampling strategies, such as neighbor sampling and layer-wise sampling, to scale the training process to large graphs.
 It also enables distributed training using PyTorch’s distributed data parallel (DDP) for faster training on multiple GPUs.
-By standardizing the GNNs, feature encoders, dataloader and evaluators, MM-Graph ensures a fair and comprehensive evaluation of graph learning algorithms on multimodal graph data.
+By standardizing the GNNs, KGEs, feature encoders, dataloader and evaluators, MM-Graph ensures a fair and comprehensive evaluation of graph learning algorithms on multimodal graph data.
 5 Empirical Analysis
-We conduct a comprehensive evaluation of our proposed benchmark on both link prediction and node classification tasks. Our experiments aim to answer the following research questions:
-(i) Are multimodal-GNNs better than conventional GNNs on MM-Graph?
-(ii) What is the best approach for encoding features? Is alignment between modalities necessary?
-(iii) How much performance gain can be achieved by using multimodal features compared to text features alone?
+We conduct a comprehensive evaluation of our proposed benchmarks on all tasks. Our experiments aim to answer the following research questions:
+(RQ1) What are the performances of MM-GNNs, GNNs, and MM-KGEs on MM-Graph?
+(RQ2) Do multimodal-GNNs consistently outperform conventional GNNs on MM-Graph?
+(RQ3) What is the most effective approach for encoding features, and is alignment between modalities necessary?
+(RQ4) How much performance gain can be achieved by using multimodal features compared to unimodal features alone?
+Figure 1: (RQ2) Multimodal GNNs underperforms conventional GNNs.
+We compare the best performance of multimodal GNNs (MMGCN/MGAT) and conventional GNNs (SAGE, GCN, BUDDY).
+Conventional GNNs consistently perform better across datasets, which justifies the importance of building MM-Graph  and calls for better multimodal GNN designs.
 To ensure fair and rigorous comparisons between various feature encoders and GNNs, we adopt consistent experimental settings across all experiments.
 We perform automatic hyperparameter tuning using Optuna [1] to optimize the performance of each model.
-Detailed experimental setup and hyperparameters are provided in the Appendix C.
-Figure 2: Feature alignment is important.
-We compare the performance of various feature encoders, find aligned features, e.g.
-CLIP and Imagebind result in much better performance compared with unaligned features on Amazon-Sports and Amazon-Cloth.
+Detailed experimental setup, hyperparameters and additional experiment results are provided in the Appendix.
+5.1 Results
+(RQ1) The detailed link prediction results are shown in Tab. 5. The detailed node classification results are shown in Tab. 7. The detailed knowledge graph completion results are shown in Tab. 6.
+Table 6: KGC results on MM-Codex-S and MM-Codex-M. Highlighted box indicates the best performing combination for each dataset. Aligned feature embeddings outperforms unaligned features.
+KGEs Dataset Image Encoder Text Encoder MRR H@1 H@3 H@10
+MoSE MM-CoDEx-s CLIP CLIP 36.59 $\pm$0.16 28.10 $\pm$0.51 41.79 $\pm$1.03 51.08 $\pm$0.29
+T5 ViT 39.00 $\pm$0.35 34.40 $\pm$0.47 42.73 $\pm$0.33 46.53 $\pm$0.49
+ImageBind ImageBind 22.45 $\pm$1.36 5.78 $\pm$1.33 35.48 $\pm$1.30 44.49 $\pm$0.98
+T5 DINOv2 40.07 $\pm$0.24 35.82 $\pm$1.85 42.77 $\pm$1.17 47.30 $\pm$0.17
+MM-CoDEx-m CLIP CLIP 6.97 $\pm$0.58 3.24 $\pm$1.11 9.33 $\pm$0.33 12.15 $\pm$0.02
+T5 ViT 6.09 $\pm$0.01 1.71 $\pm$0.22 8.42 $\pm$0.26 12.56 $\pm$0.28
+ImageBind ImageBind 7.01 $\pm$0.15 2.26 $\pm$0.33 9.60 $\pm$0.26 13.98 $\pm$0.24
+T5 DINOv2 6.40 $\pm$0.18 2.14 $\pm$0.22 9.06 $\pm$0.23 12.74 $\pm$0.14
+VISTA MM-CoDEx-s CLIP CLIP 28.49 $\pm$0.39 19.12 $\pm$0.64 30.53 $\pm$0.46 49.83 $\pm$0.31
+T5 ViT 29.70 $\pm$0.51 20.00 $\pm$0.59 32.11 $\pm$0.31 50.52 $\pm$0.75
+ImageBind ImageBind 30.39 $\pm$0.65 20.20 $\pm$0.84 33.52 $\pm$0.30 52.68 $\pm$0.50
+T5 DINOv2 26.98 $\pm$1.68 17.81 $\pm$1.16 29.53 $\pm$2.58 47.28 $\pm$2.86
+MM-CoDEx-m CLIP CLIP 22.19 $\pm$1.92 15.81 $\pm$1.62 24.31 $\pm$2.10 34.73 $\pm$2.58
+T5 ViT 22.10 $\pm$1.61 15.92 $\pm$1.33 24.09 $\pm$1.71 34.68 $\pm$2.26
+ImageBind ImageBind 23.20 $\pm$1.45 16.97 $\pm$1.35 24.95 $\pm$1.24 35.88 $\pm$1.91
+T5 DINOv2 21.38 $\pm$1.17 15.35 $\pm$1.15 23.19 $\pm$1.06 33.40 $\pm$1.49
+Here are our findings:
+(RQ2) Multimodal GNNs perform poorly, while GNNs and MM-KGES often outperforms on the proposed MM-Graph datasets.
+Our experiments, as illustrated in Figure 1 and Tab.
+5, Tab.
+7, and Tab.
+6, reveal an unexpected trend: multimodal GNNs specifically designed for processing multimodal input data, such as MMGCN and MGAT, do not consistently outperform conventional GNN models like SAGE across various tasks and datasets.
+This counterintuitive result may be attributed to the architectural designs of MMGCN and MGAT, which primarily perform message passing and aggregation on each modality separately, only fusing information across modalities at the final stage.
+We hypothesize that this approach leads to insufficient integration or alignment of multimodal information, highlighting the need for more sophisticated model designs for multimodal graphs.
+Figure 2: (RQ3) Feature alignment is important.
+We compare the performance of various feature encoders, find aligned features, e.g., CLIP and Imagebind result in much better performance compared with unaligned features on Amazon-Sports and Amazon-Cloth.
 Among them, Imagebind performs the best across backbones.
 This indicates the importance of using aligned features on these datasets.
-5.1 Results
-The detailed link prediction results on Amazon-Sports, Amazon-Cloth and Goodreads-LP is shown in Table  3. The detailed node classification results on is shown in Table  4. Here are our findings:
-Multimodal GNNs perform poorly on the proposed MM-Graph datasets.
-As shown in Fig  5.1, we observe that MMGCN and MGAT, the two multimodal GNNs specifically designed to process multimodal input data does not perform better than conventional GNN models such as SAGE.
-As MMGCN and MGAT mostly just perform message passing and aggregation on each modality separately and then fuse the information across modality together at last, we hypothesize that it is likely to do the insufficient fusion of modalities, and calls for better model designs of multimodal structured data.
-Feature Encoders (Alignment).
-CLIP and ImageBind are two feature encoders which can map features of various modalities to the same embedding space while T5ViT and T5Dinov2 uses SOTA embedding methods for each modality separately, without any specific alignment layers.
-Interesting, we found the results to vary across datasets.
-As shown in Figure  2, we found that for Amazon-Sports and Amazon-Cloth, the two amazon datasets, aligned multimodal features performs consistently better than unaligned features.
-Specifically, we also found that ImageBind consistently gives the best results across backbones.
-On Goodreads-LP, we found that T5Dinov2 performs best most of the times, this indicates that Dinov2 excels other vision models in the Optical Character Recoginition (OCR) task.
-Table 4: Node classification results on Goodreads-NC and Ele-Fashion. *OOM = out of GPU memory.
+In contrast, we observed that MoSE and VISTA perform notably well on the two proposed MMKG datasets, with the exception of MoSE on MM-CoDEx-m.
+MoSE’s approach of learning separate KGE embeddings for each modality and then ensembling them together represents a form of late fusion.
+While this method shows promise, its performance variability across datasets suggests that the effectiveness of late fusion strategies may be dataset-dependent.
+Our new findings underscore the complexity of effectively leveraging multimodal information in graph learning tasks.
+They also highlight the need for further research into more advanced fusion techniques that can better capture and integrate the complementary information provided by different modalities throughout the learning process, rather than relying on late-stage integration.
+(RQ3) Feature alignment is important for multimodal graph learning.
+Our evaluation of four feature encoders, CLIP, ImageBind, T5 + ViT, and T5 + DINOv2, reveals significant performance variations across datasets and tasks.
+CLIP and ImageBind, which map features from various modalities to a shared embedding space, consistently outperform T5 + ViT and T5 + DINOv2, which employ independent embedding methods for each modality without specific alignment layers.
+Table 7: Node classification results on Goodreads-NC and Ele-Fashion.
+Conventional GNNs such as SAGE performs best across datasets.
+Aligned feature embeddings, e.g.
+CLIP and ImageBind outperforms unaligned features.
 Image Encoder Text Encoder Ele-fashion Goodreads-NC
-MMGCN CLIP CLIP 86.10 $\pm$0.50 OOM
-T5 ViT 82.39 $\pm$0.30 OOM
-ImageBind ImageBind 86.21 $\pm$0.94 OOM
-T5 Dinov2 85.53 $\pm$0.33 OOM
-MGAT CLIP CLIP 84.66 $\pm$0.29 OOM
-T5 ViT 84.01 $\pm$0.08 OOM
-ImageBind ImageBind 86.12 $\pm$0.08 OOM
-T5 Dinov2 84.54 $\pm$0.27 OOM
+MMGCN CLIP CLIP 86.10 $\pm$0.50 83.29 $\pm$0.20
+T5 ViT 82.39 $\pm$0.30 81.85 $\pm$0.22
+ImageBind ImageBind 86.21 $\pm$0.94 80.58 $\pm$1.08
+T5 DINOv2 85.53 $\pm$0.33 82.44 $\pm$0.11
+MGAT CLIP CLIP 84.66 $\pm$0.29 76.48 $\pm$0.59
+T5 ViT 84.01 $\pm$0.08 75.43 $\pm$0.76
+ImageBind ImageBind 86.12 $\pm$0.08 69.45 $\pm$6.25
+T5 DINOv2 84.54 $\pm$0.27 74.98 $\pm$1.23
 GCN CLIP CLIP 79.83 $\pm$0.03 81.61 $\pm$0.01
@@
 T5 DINOv2 84.87 $\pm$0.01 68.83 $\pm$0.03
-Figure 3: Multimodal GNNs underperforms conventional GNNs.
-We compare the best performance of multimodal GNNs (MMGCN/MGAT) and conventional GNNs (SAGE, GCN, BUDDY).
-Conventional GNNs consistently perform better across datasets, which justifies the importance of building MM-Graph  and calls for better multimodal GNN designs.
-[…]
-5.2 Challenges and Future Directions.
-First, existing multimodal GNNs are all based on full-batch training schemes and it would go out of GPU memory when the graph is large, e.g.
-Goodreads-NC and Goodreads-LP.
-Mini-batch multimodal GNNs are worth exploring in the future.
-Second, we find that it is hard to find a graph dataset with audio/video features with open-source access, even though music/video recommendation is an task of great importance in real-world deployed systems.
-We believe the study is similar to what is conducted here and the ImageBind features we use can be easily extended to these scenarios.
-However, integrating multimodal graphs with audio/video features still require future exploration.
-Figure 4: Multimodal features are helpful for graph learning.
-Multimodal features, performs better than text features across datasets and tasks, which justifies the necessity of introducing multimodal graph datasets.
+Figure 3: (RQ4) Multimodal features are helpful for graph learning.
+Multimodal features perform better than text-only features across datasets and tasks, which justifies the necessity of introducing multimodal graph datasets.
+For link prediction, aligned multimodal features consistently outperform unaligned features, as illustrated in Figure 2.
+ImageBind yields the best results across all backbones.
+However, on the Goodreads-LP dataset, T5+DINOv2 demonstrates superior performance due to its robust Optical Character Recognition (OCR) capabilities, which are crucial for extracting textual information from book covers.
+In node classification tasks, ImageBind and CLIP again emerge as the top-performing feature encoders.
+This further emphasizes the importance of feature alignment in graph-based tasks.
+Similarly, for knowledge graph completion, ImageBind consistently demonstrates superior performance across datasets and various KGEs, except for MM-CoDEx-s on MoSE.
+For MM-CoDEx-s on MoSE, we suspect that the modality split function in MoSE leads to unaligned embedding space for KGEs and thus hurts the performance.
+On Goodreads-LP, the T5 + DINOv2 combination demonstrates superior performance.
+We believe it is because the Goodreads dataset presents a distinctive challenge because book covers primarily contain textual information, requiring robust capabilities of extracting and understanding text from book covers.
+This is crucial because author information is important when predicting reading preferences, and it is only displayed on book covers.
+Users tend to prefer books by the same author, making OCR-capable models particularly effective in this context.
+Our findings align with Tong et al.
+[40], where CLIP-based VLM is shown to be suboptimal at capturing details.
+(RQ4) Multimodal models outperform unimodal models, underscoring the importance of visual information.
+To highlight the necessity of multimodal graph benchmarks beyond existing text-rich graphs, we conducted a study comparing the performance gains achieved by multimodal features over unimodal text or visual features alone (T5/ViT).
+Using an MLP model with optimal hyperparameters selected through Optuna, we focused on isolating the effects of feature types.
+As shown in Figure 3, integrating multimodal features consistently results in more than a 6% improvement in performance across various datasets and tasks.
+While visual information alone does not surpass text in performance, the combination of both modalities significantly outperforms any single modality.
+This finding underscores the critical role of visual information in enhancing graph representations.
+Note that we do not report results for knowledge graph completion (KGC) tasks, as all evaluated KGEs require multimodal inputs.
 6 Conclusion
-In summary, this paper introduces the Multimodal Graph Benchmark (MM-Graph), a comprehensive benchmark for multimodal graph data.
-MM-Graph consists of five diverse datasets spanning various domains and tasks, featuring rich textual and visual node features.
-To enable a fair and extensive evaluation, MM-Graph standardizes the GNN architectures, feature encoders, dataloaders and evaluators used across all datasets.
-The GNNs include both unimodal and multimodal architectures, while the feature encoders leverage state-of-the-art models for capturing semantic information from text and images.
-Experiments on MM-Graph datasets provide valuable insights into the performance of different GNN and feature encoder combinations in multimodal graph learning settings.
-By offering a unified evaluation framework, MM-Graph aims to facilitate research on multimodal graph representation learning and drive the development of more powerful and robust graph learning algorithms for real-world applications.
-Social impacts.
-MM-Graph has the potential to significantly advance research on multimodal graph representation learning, which can enable novel applications that positively transform various aspects of society.
-Many real-world domains, such as healthcare, social networks, and urban computing, involve multimodal graph data.
-By providing a comprehensive evaluation framework, MM-Graph can accelerate the development of more powerful graph learning algorithms for these domains.
-MM-Graph also promotes reproducible research by offering standardized benchmarks, fostering collaboration within the research community.
-Ultimately, MM-Graph aims to unlock the value of multimodal graph data to enable impactful real-world applications.
-However, as with all technologies of this nature, it is important to consider how it can be used responsibly and with responsible development in mind.
-There are risks of these algorithms perpetuating or amplifying societal biases if the underlying graph data contains biased patterns.
-Additionally, misuse of multimodal graph can cause disinformation and targeted manipulation.
-It is therefore crucial that researchers using MM-GRAPH prioritize the responsible development of multimodal graph algorithms in line with ethical principles.
+This paper presents MM-Graph, a comprehensive benchmark for multimodal graph learning.
+MM-Graph includes seven diverse datasets with rich textual and visual features, standardizing GNN architectures, KGEs, feature encoders, dataloaders, and evaluators to ensure fair evaluation.
+This setup provides insights into the performance of various GNN and encoder combinations in multimodal settings.
+While MM-Graph advances the integration of text and visual data, challenges remain in incorporating audio and video features due to the scarcity of open-source datasets.
+Although ImageBind features offer a basis for future research, further exploration is needed to effectively integrate these modalities.
+MM-Graph aims to drive the development of robust graph learning algorithms for real-world applications by highlighting the importance of visual information and setting the stage for future innovations in multimodal graph learning.
+Acknowledgments
+This material is based upon work supported by a SnapChat gift, the National Science Foundation under IIS 2212143 and CAREER Grant No.
+IIS 1845491.
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or other funding parties.
 Appendix A Graph Statistics Metrics
@@
 These metrics enable researchers to assess the strengths and weaknesses of different graph learning algorithms in ranking positive edges higher than negative edges, a crucial task in many real-world applications involving graph data.
-Appendix C Experimental details
+Appendix C Other Candidate Datasets
+We also provide a list of candidate datasets and the reasons why they are unsuitable for the our benchmark here.
+Ogbn-Arxiv: Raw paper IDs are not provided and cannot be retrieved by backtracking, making it impossible to crawl images from Arxiv.
+Yelp: Only reviews of restaurants have textual and visual information.
+However, the graph structure of user-friend-with-user exists, and typically, a user’s friends do not have any reviews on any restaurant.
+Additionally, Yelp’s license prevents repurposing the datasets.
+Tiktok: This dataset does not provide any raw text or image features.
+Appendix D Experimental details
 To ensure a fair and comprehensive evaluation of different graph learning algorithms on the Multimodal Graph Benchmark (MM-Graph), we conduct extensive experiments with rigorous experimental settings and hyperparameter tuning.
 In this section, we provide detailed information about the experimental setup, hyperparameter search process, and optimization strategies employed in our study.
-Table 5: Best performing hyperparameters under Optuna search on Amazon-Sports, Amazon-Cloth and Goodreads-LP. *OOM = out of GPU memory.
-Encoder Amazon-Sports Amazon-Cloth Goodreads-LP
-Image Text LR # Layers LR # Layers LR # Layers
-MMGCN CLIP 0.002595055216 1 0.00120638677883 2 OOM OOM
-ViT T5 0.00372938084 1 0.00153969042325 2 OOM OOM
-ImageBind 0.0008025598304 1 0.00051054426806 1 OOM OOM
-Dinov2 T5 0.001184411766 2 0.00157538063161 2 OOM OOM
-MGAT CLIP 0.001706100258 1 0.000109845035707 3 OOM OOM
-ViT T5 0.001673818621 3 0.000242709240618 3 OOM OOM
-ImageBind 0.000221326243 2 0.0001492753647 2 OOM OOM
-Dinov2 T5 0.0001216520274 2 0.00007883247879 3 OOM OOM
-GCN CLIP 0.00199258857 1 0.001250668422 1 0.0004599971993 1
-ViT T5 0.0004748389636 1 0.0003867480145 1 0.0002695086535 1
-ImageBind 0.0002173991668 1 0.00025841621630 1 0.0003440800285 1
-Dinov2 T5 0.0003867480145 1 0.00050429857558 1 0.0005790300826 2
-SAGE CLIP 0.000888801759 3 0.0006242779054 3 0.001115782634 3
-ViT T5 0.0009069080338 3 0.0009159332036 3 0.001115782634 3
-ImageBind 0.0007536062461 3 0.001115782634 3 0.0004026026641 3
-Dinov2 T5 0.0009159332036 3 0.0009159332036 3 0.001115782634 3
-BUDDY CLIP 0.0001996185501 1 0.0001073125661 2 0.0003131657128 1
-ViT T5 0.0001782348253 1 0.000104734288 1 0.000101783244 1
-ImageBind 0.0001082982111 3 0.0001048823969 3 0.001856957299 1
-Dinov2 T5 0.0002695086535 1 0.0001385185507 1 0.0001996185501 1
-MLP CLIP 0.0005982646116 1 0.0004259767004 1 0.0002317306399 3
-ViT T5 0.00027681242 3 0.0003867480145 1 0.0001156508232 3
-ImageBind 0.0001511847491 2 0.0001730467441 2 0.000101783244 3
-Dinov2 T5 0.0002317306399 3 0.0003867480145 1 0.0001156508232 3
-Table 6: Best performing hyperparameters under Optuna search on Ele-fashion and Goodreads-NC. *OOM = out of GPU memory.
-Encoder Ele-fashion Goodreads-NC
-Image Text LR # Layers LR # Layers
-MMGCN CLIP 0.01617331603 1 OOM OOM
-ViT T5 0.0001215688562 2 OOM OOM
-ImageBind 0.01590607512 2 OOM OOM
-Dinov2 T5 0.01233773648 1 OOM OOM
-MGAT CLIP 0.0003211703387 1 OOM OOM
-ViT T5 0.0004549756184 2 OOM OOM
-ImageBind 0.003739550104 2 OOM OOM
-Dinov2 T5 0.0002344834419 1 OOM OOM
-GCN CLIP 0.002747501509 2 0.0004599971993 1
-ViT T5 0.0007123650663 2 0.0002695086535 1
-ImageBind 0.0007123650663 2 0.0003440800285 1
-Dinov2 T5 0.0007123650663 2 0.0005790300826 2
-SAGE CLIP 0.003235718844 3 0.001115782634 3
-ViT T5 0.001523236953 2 0.001115782634 3
-ImageBind 0.00174700527 2 0.0004026026641 3
-Dinov2 T5 0.00134217448 2 0.001115782634 3
-MLP CLIP 0.004407317752 3 0.00174700527 3
-ViT T5 0.001342211088 2 0.0004414532347 3
-ImageBind 0.00278539483 2 0.0001409760217 3
-Dinov2 T5 0.001287758261 3 0.0004611586099 3
-C.1 Hyperparameter Tuning
+D.1 Hyperparameter Tuning
 Proper hyperparameter tuning is crucial for obtaining reliable and meaningful results when evaluating machine learning models.
@@
 Each experiment is run three times with different random seeds, and we report the mean and standard deviation of the respective performance metrics.
-C.2 Optimization
+D.2 Optimization
 All experiments are run with 3 different random seeds to account for the stochasticity in model training and initialization.
@@
 This allows the model to make larger updates in the early stages of training and fine-tune the parameters in the later stages.
-The best-performing hyperparameters for each combination of feature encoders and GNNs are presented in Table 5 and 6.
-These hyperparameters are selected based on the highest mean performance across the 3 runs.
-C.3 Reproducibility
-We make the source code for MM-Graph publicly available on GitHub, along with detailed documentation and instructions for running the experiments.
-The datasets used in MM-Graph are also made accessible through the Hugging Face Datasets platform, allowing researchers to easily access and use the data for their own experiments.
-By providing a comprehensive experimental setup, detailed hyperparameter configurations, and open-source code and datasets, we aim to foster reproducibility and encourage further research on multimodal graph learning using the MM-Graph benchmark.
-Appendix D Dataset
-In this section, we provide details on the release and maintenance of the Multimodal Graph Benchmark (MM-Graph) datasets, the file structure of the datasets, and instructions on how to use them effectively.
-D.1 Dataset Release
-To facilitate easy access and widespread use of the MM-Graph datasets, we have created a dedicated webpage at https://mm-graph-benchmark.github.io/.
-The webpage provides a direct link to the datasets and showcases representative examples from each dataset, allowing researchers to quickly grasp the nature and diversity of the data.
-The MM-Graph datasets are hosted on the Hugging Face Datasets platform at https://huggingface.co/datasets/mm-graph-org/mm-graph.
-Hugging Face Datasets is a popular and reliable platform for hosting and sharing datasets in the machine learning community.
-By leveraging this platform, we ensure the long-term maintenance and preservation of the MM-Graph datasets.
-We plan to periodically update the repository to address any issues or incorporate feedback from the community.
-The MM-Graph datasets are released under the Apache 2.0 license, which allows for open access and permissive use of the data for research purposes.
-This licensing choice aligns with our goal of fostering collaborative research and encouraging the development of advanced multimodal graph learning algorithms.
-To complement the datasets, we also provide a standardized benchmark, evaluator, and various GNN backbones and feature encoders on our GitHub repository at https://github.com/mm-graph-benchmark/mm-graph-benchmark.
-The repository includes detailed documentation and example usage to facilitate the adoption and extension of the MM-Graph benchmark.
-D.2 Dataset File Structure
-The MM-Graph directory contains five separate folders, each corresponding to a specific dataset: amazon-sports, amazon-cloth, goodreads-lp, goodreads-nc, and ele-fashion.
-Each folder contains the necessary files for the respective dataset.
-For the three link prediction datasets (amazon-sports, amazon-cloth, and goodreads-lp), each folder contains the following files:
-  • lp-edge-split.pt: This file contains the graph structures, along with the train, validation, and test edges, as well as their corresponding negative pairs.
-  • clip_feat.pt, t5vit_feat.pt, imagebind_feat.pt, t5dino_feat.pt: These files contain the node features obtained from various feature encoders, as discussed in Section 4.2 of the paper.
-For the two node classification datasets (goodreads-nc and ele-fashion), each folder contains the following files:
-  • edges.pt: This file contains the graph structures.
-  • split.pt: This file indicates the train, validation, and test nodes.
-  • label-w-missing.pt: This file stores the node labels. Nodes with missing labels are assigned to a dummy class and are not used as train, validation, or test nodes.
-  • clip_feat.pt, t5vit_feat.pt, imagebind_feat.pt, t5dino_feat.pt: These files contain the node features obtained from various feature encoders, as discussed in Section 4.2 of the paper.
-D.3 Dataset Usage
-To load and use the MM-Graph datasets, we provide a user-friendly Python package.
-For detailed examples and usage instructions, please refer to the GitHub repository.
-The repository provides comprehensive documentation, including installation instructions, dataset descriptions, and example scripts for training and evaluating models on the MM-Graph datasets.
-By providing a standardized and easy-to-use interface for accessing the datasets, along with detailed documentation and examples, we aim to facilitate the adoption of the MM-Graph benchmark and enable researchers to focus on developing and evaluating their multimodal graph learning models.
+D.3 Additional Results
+SOTA methods applied to text-only data.
+We compare ImageBind+SAGE with UniGraph, one of the SOTA methods on text-attributed graphs, and results shown in Tab.
+8 are from  [14].
+It demonstrates that multimodal inputs consistently outperform text-only inputs.
+Table 8: Node classification results on Goodreads-NC and Ele-Fashion.
+Dataset UniGraph ImageBind+SAGE
+Amazon-Sports 27.11 $\pm$ 0.10 34.32 $\pm$ 0.11
+Amazon-Cloth 18.01 $\pm$ 0.03 25.20 $\pm$ 0.09
+Goodreads-LP 22.31 $\pm$ 0.05 34.61 $\pm$ 0.43
+Goodreads-NC 78.14 $\pm$ 0.11 80.39 $\pm$ 0.21
+Ele-fashion 81.05 $\pm$ 0.08 87.71 $\pm$ 0.13
+F1 score
+Additionally, we report F1-score on the two node classification datasets. The results are shown below.
+Table 9: Additional results on Macro-F1 score
+Image Encoder Text Encoder Ele-fashion Goodreads-NC
+MMGCN CLIP CLIP 75.07 $\pm$0.45 81.02 $\pm$0.32
+T5 ViT 63.98 $\pm$1.18 79.45 $\pm$0.49
+ImageBind ImageBind 76.18 $\pm$0.72 78.41 $\pm$0.45
+T5 DINOv2 73.62 $\pm$0.04 80.13 $\pm$0.13
+GCN CLIP CLIP 79.85 $\pm$0.04 74.36 $\pm$0.01
+T5 ViT 79.69 $\pm$0.08 76.27 $\pm$0.03
+ImageBind ImageBind 80.47 $\pm$0.12 72.84 $\pm$0.10
+T5 DINOv2 79.39 $\pm$0.09 81.72 $\pm$0.08
+SAGE CLIP CLIP 87.08 $\pm$0.01 83.30 $\pm$0.04
+T5 ViT 85.81 $\pm$0.04 82.94 $\pm$0.08
+ImageBind ImageBind 87.72 $\pm$0.08 80.26 $\pm$0.04
+T5 DINOv2 85.45 $\pm$0.12 82.88 $\pm$0.02